### v1

# Boundary Attention: Learning to Find Faint Boundaries at Any Resolution

Abstract
We present a differentiable model that explicitly models boundaries—including contours, corners and junctions—using a new mechanism that we call boundary attention. We show that our model provides accurate results even when the boundary signal is very weak or is swamped by noise. Compared to previous classical methods for finding faint boundaries, our model has the advantages of being differentiable; being scalable to larger images; and automatically adapting to an appropriate level of geometric detail in each part of an image. Compared to previous deep methods for finding boundaries via end-to-end training, it has the advantages of providing sub-pixel precision, being more resilient to noise, and being able to process any image at its native resolution and aspect ratio.
1 Introduction
Converting a precise contour that is defined mathematically in continuous 2D space to its discrete representation in pixel-space is a common task in computer graphics, and there are established techniques for rasterization [12, 27], anti-aliasing [16, 6] and so on. However, the inverse problem in computer vision of robustly inferring precise and unrasterized descriptions of contours from discrete images remains an open challenge. Earlier work in the field inferred unrasterized representations of edges [4, 8, 25, 14, 13, 22]—and sometimes corners and junctions too—and it explored algorithms for connecting them into boundaries via edge-linking and MRFs. But since the dawn of deep learning, boundaries have almost exclusively been modeled using discrete, rasterized maps. Moreover, it is not uncommon to rely heavily on supervision from human annotations, and even modern datasets can suffer from annotation errors on the order of pixels, quite far from sub-pixel accuracy.
Taking inspiration from early computer vision work, we present a model for inferring unrasterized boundaries that can be learned instead of designed. We show that this model is able to benefit from the power of deep learning while also realizing many of the advantages—such as robustness to noise, sub-pixel precision, and adaptability between signal-types—that have long been the potential strengths of classical bottom-up techniques.
The core of our model is a mechanism we call boundary attention. It is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specifies the local boundaries surrounding every pixel. The model’s output is a field of overlapping geometric primitives that can be used in a variety of ways, including to produce an unsigned distance function for the image’s boundaries, a boundary-aware smoothing of its channel-values, and a field of soft local attention maps that relate every pixel to its neighbors.
To evaluate our model, we consider the problem of finding boundaries in images corrupted by extreme amounts of noise [23, 24, 33], where the images can be of any size and resolution, and where we know little about the noise and the objects or shapes we are looking for. We choose this problem because it relies entirely on having a strong model for the basic topological and geometric properties of boundaries, namely that they are piecewise smooth curves which connect at corners or junctions and divide regions of homogeneity.
Figure 1: Our model uses neighborhood attention with dense, stride-1 tokens (but we depict non-overlapping ones here for visual clarity). The entire model is invariant to discrete spatial shifts and so applies at any resolution. Each token encodes an adaptively-sized geometric primitive that represents the unrasterized local boundaries surrounding a pixel, and the tokens evolve through boundary attention to be geometrically consistent. The output is an overlapping field of primitives that directly implies a boundary-aware smoothing of the input image and an unsigned distance map to the image boundaries.
We show that our model has several advantages. All of its components are local and invariant to discrete spatial shifts, so it can be trained on small-sized images and then deployed on much larger and differently-shaped ones. It is also very compact, comprising only 207k parameters, and it runs several times faster than many alternatives. We also find that it can be trained to a useful state with very simple synthetic data, made up of random circles and triangles that are uniformly colored and then corrupted by noise. Despite this simplicity, we find that the model can generalize to real images surprisingly well.
Our main contributions can be summarized as follows:
  1. 1.
    We propose a novel network design that explicitly models boundaries using a boundary attention mechanism that can be used in any deep learning framework.
  2. 2.
    We demonstrate that our model based on boundary attention is more effective than current state-of-the-art methods at finding boundaries at high levels of sensor noise (Fig. LABEL:fig:real_results) and can achieve sub-pixel accuracy, even in the presence of noise.
2 Related Work
It is possible to process the outputs of certain filters to localize edges with sub-pixel precision (e.g.,  [4, 8, 25, 14, 13, 22]). But this approach struggles near corners and junctions because the edge filters are not derived for these, and accuracy breaks down. Recently, the field of junctions [33] showed improvements by expanding the library of local, unrasterized geometric primitives from classical edge-based primitives to a larger family of “generalized junctions”. This allows unifying a variety of bottom-up cues that had often previously been handled separately, such as geometrically-consistent and polarity-consistent edge elements being linked into contours [18, 23, 24, 4]; contours that approach one another being joined at corners and junctions [15, 21, 36]; homogeneous regions that look the same being together; and those that look different being apart [7, 11, 2]. The field of junctions showed that putting all of these cues together leads to a dramatic improved in noise resilience, and it also allows analyzing the recovered boundaries explicitly into its component junctions, edges and corners. In this paper, we leverage all of these benefits, and we do this more efficiently and with an adaptive patch size. Additionally, our model is differentiable, meaning it can be used as a component in larger systems where accurate boundaries are necessary.
Our work is complementary to recent approaches that leverage large-scale training to perform edge detection or segmentation by internalizing the patterns in a training set. Recent examples include Segment Anything [19] and combinations of strong self-supervised features with clustering and CRFs [5, 3, 20]. These large-scale approaches rely on large training sets and high-capacity architectures that can internalize high-level cues like object and scene familiarity as well as low-level cues like spatial coherence. Our approach is different because it relies exclusively on low-level cues, has much less capacity, and uses much less training data. Despite this difference, we find that our model’s inherent preference for contour-like boundaries allows it to achieve comparable (or better) results, in particular when noise is high. This suggests that future work could benefit from incorporating some of our architectural elements into larger and higher-capacity systems, in particular to improve interpretability, efficiency, generalization or task-transfer.
3 Representation
Our system is depicted in Figure 1. It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with pixel-resolution tokens. Critically, each pixel’s token is explicitly required to encode a tuple of values representing a geometric primitive that specifies the geometric structure of the local boundaries surrounding the pixel. Because the tokens are forced through this geometric bottleneck, we refer to our model’s core mechanism as Boundary Attention.
We instantiate our bottleneck using a learned linear mapping from the token dimension to a predefined lower-dimensional space of unrasterized boundary patterns that we call junction space. Our choice for this space is inspired by the Field of Junctions [33], but we use a modified parameterization (described below) that is differentiable across its entire domain. Junction space has the benefit of specifying local boundary patterns without rasterization and thus with unlimited spatial precision. As depicted in Figure 2 and described in [33], it also has the benefit of including a large family of local boundary patterns, including uniformity (i.e., absence of boundaries), edges, bars, corners, T-junctions and Y-junctions.
The next section introduces our parameterization of junction space and some associated operators. Then Section 4 describes the architecture that we use to analyze an image into its field of junction values.
Figure 2: Samples on a smooth trajectory through our space of geometric primitives (“junction space”), with one sample enlarged and accompanied by a visualization of its distance map on the right. Each sample comprises a vertex position $(u,v)$, orientation $\theta$, and angles $(\omega_{1},\omega_{2},\omega_{3})$ that partition the circle. The space includes all types of locally-uniform regions, edges, bars, corners, Y-junctions and T-junctions.
3.1 Boundary Primitives
Throughout this paper we use parentheses $(x)$ for continuous signals defined on the 2D image plane $[0,W]\times[0,H]$ and square brackets $[n]$ for discrete signals defined on the pixel grid. We use $c[n]$ for the coordinates of the $n$th pixel’s center.
Denote the $K$-channel input image by $\{\mathbf{f}[n]\}$, where $\mathbf{f}[n]\in\mathbb{R}^{K}$ is the vector image value at the discrete pixel grid index $n$. Our approach is to analyze the image into a field of dense, stride-1 overlapping local patches, each having a square support $\Omega_{n}(x)$ centered at the $n$th pixel.
There are many ways to partition a local region $\Omega_{n}(x)$, and one can define parametric families of such partitions. For example the set of oriented lines provides a two-parameter family of partitions, with each member of the family separating the region into points that lie on one side of a line or the other. This family of partitions would be appropriate for describing edges. Here we define a larger family of partitions that encompasses a greater variety of local boundary structures.
As depicted in the right of Figure 2, our partitions are parameterized by $\mathbf{g}\in\mathbb{R}^{2}\times\mathbb{S}^{1}\times\triangle^{2}$, where $\mathbb{S}^{1}$ is the unit circle and $\triangle^{2}$ is the standard $2$-simplex. We use the notation $\mathbf{g}=(\boldsymbol{u},\theta,\boldsymbol{\omega})$, where $\boldsymbol{u}=(u,v)\in\mathbb{R}^{2}$ is the vertex, $\theta\in\mathbb{S}^{1}$ is the orientation, and $\boldsymbol{\omega}=(\omega_{1},\omega_{2},\omega_{3})$ are barycentric coordinates (defined up to scale) for the three relative angles, ordered clockwise starting from $\theta$. Our convention is to express the vertex coordinates relative to the center of region $\Omega_{n}(x)$, and we denote that the vertex is free to move outside of this region. We also note that up to two angles $\omega_{j}$ can be zero. This all makes it possible to represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions).
Fixing a value for $\mathbf{g}$ induces three binary-valued spatial support functions:
$s_{nj}(x;\mathbf{g})\in\{0,1\},\ j=1,2,3,$ (1)
that evaluate to $1$ for points $x$ that are in $\Omega_{n}(x)$ and in the $j$th wedge defined by $\mathbf{g}$; and 00 otherwise. It also induces an unsigned distance function:
$d_{n}(x;\mathbf{g})\geq 0,$ (2)
that represents the Euclidean distance from point $x$ to the nearest point in the boundary set defined by $\mathbf{g}$.
Defining these as continuous functions allows specifying the local boundary structure with unlimited spatial resolution. The right of Figure 2 uses three colors to visualise the wedge supports of a junction $\mathbf{g}$, and it shows a quantized visualization of the associated distance function.
Instead of fixing the size of every patch in the image to a hand-selected value that is the same throughout, we would like the size of every region $\Omega_{n}$ to adapt to the local geometry. We do this by equipping each patch with a parameterized local windowing function $w_{n}(x;\mathbf{p})\in[0,1]$. Specifically, we parameterize them as convex combinations of $W$ square window function coefficients $\mathbf{p}\in{\cal P}=\triangle^{W-1}$. That is,
$w_{n}(x;\mathbf{p})=\sum_{i=1}^{W}p_{i}\mathbf{1}[\|x-c[n]\|_{\infty}\leq D_{i}],$ (3)
where $\|\cdot\|_{\infty}$ is the $\ell^{\infty}$-norm, and $\mathbf{1}[\cdot]$ is the indicator function that returns $1$ if the argument true; and 00 otherwise. In our experiments we use $W=3$ and diameters $\mathbf{D}=(3,9,17)$.
Figure 3: Visualizing our model’s output. Top row, left to right: (i) input image; (ii) boundary-aware smoothing; and (iii) boundary map with attention maps at four query points. The two bottom rows show regions $\Omega_{n}$ unfolded from within the two white windows indicated above. Bottom rows, left to right: (i) window functions; (ii) boundaries; and (iii) support functions colored by their gathered wedge features. See text for details.
Figure 4: Model Architecture. All blocks are invariant to discrete spatial shifts, and only colored blocks are learned. Orange blocks operate at individual locations $n$, while blue ones operate on small spatial neighborhoods. Symbol $\oplus$ is concatenation, and gather and slice operators (Eqs. 46) are depicted at right. The first iteration uses $\boldsymbol{\gamma}^{0}[n]=\boldsymbol{\gamma}_{0}[n]$, $\bar{\mathbf{f}}^{0}[n]=\mathbf{f}[n]$, and $\boldsymbol{\pi}^{0}[n]=\boldsymbol{\pi}_{o}$ with $\boldsymbol{\pi}_{o}$ learned across the training set. Boundary attention repeats $T=8$ times, with one set of weights for the first four iterations and another for the last four iterations. There are $207$k trainable parameters in total.
3.2 Gather and Slice Operators
Our network operates by refining the field $\{(\mathbf{g}^{t}[n],\mathbf{p}^{t}[n])\}$ over a fixed sequence of steps $t=1,...,T$. It uses two operators that we define here and depict in the right of Figure 4. The first operator is a patch-wise gather operator, in which each wedge of each region computes the weighted average of the image values it contains (recall that $c[n]$ are $n$th pixel’s coordinates):
$\mathbf{f}_{kj}=\frac{\sum_{n}\mathbf{f}[n]w_{k}(c[n];\mathbf{p}[n])s_{kj}(c[n];\mathbf{g}[n])}{\sum_{n}w_{k}(c[n];\mathbf{p}[n])s_{kj}(c[n];\mathbf{g}[n])}.$ (4)
The second operation is a pixel-wise slice operation, where each pixel computes the means and variances, over all regions that contain it, of the per-region distance maps $d_{n}(x;\mathbf{g}[n])$ and gathered wedge features $\mathbf{f}_{kj}$. The expressions for the means are:
$\bar{d}[n]$ $=\frac{\sum_{k}w_{k}(c[n];\mathbf{p}[k])d_{k}(c[n];\mathbf{g}[k])}{\sum_{k}w_{k}(c[n];\mathbf{p}[k])},$ (5)
$\bar{\mathbf{f}}[n]$ $=\frac{\sum_{k}w_{k}(c[n];\mathbf{p}[k])\sum_{j}\mathbf{f}_{kj}s_{kj}(c[n];\mathbf{g}[k])}{\sum_{k}w_{k}(c[n];\mathbf{p}[k])\sum_{j}s_{kj}(c[n];\mathbf{g}[k])}.$ (6)
Note that the only contributions to the sums over patches $k$ are from patches that contain $c[n]$, i.e., the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$. The expressions for the pixel-wise distance map variance $\nu_{d}[n]$, and for the feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are defined accordingly and included in the supplement.
3.3 Visualizing Output
At our network’s output, we expect the shapes of junction boundaries in overlapping regions to agree, so that the variances $\nu_{d}[n],\nu_{f}[n]$ are small at every pixel. Then, the fields of means $\{\bar{d}[n]\},\{\bar{\mathbf{f}}[n]\}$ can be interpreted, respectively, as a global unsigned distance map for the image boundaries and a boundary-aware smoothing of its input channel values. Figure 3 shows an example, where we visualize the zero-set of the global unsigned distance map—we call this the global boundary map—by applying the nonlinearity:
$\bar{b}_{\eta}[n]=\left(1+(\bar{d}[n]/\eta)^{2}\right)^{-1},$ (7)
setting $\eta=0.7$.
The output contains much more information than these fields of means. For any query pixel $n$, we can probe the wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ that contain it, thereby obtaining a spatial attention map $a_{n}(x)$ that surrounds the query pixel. This is the boundary-aware spatial kernel that turns a neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the value $\bar{\mathbf{f}}[n]$:
$\bar{\mathbf{f}}[n]=\sum_{k}a_{n}(c[k])\mathbf{f}[k].$ (8)
The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6, and its maximum diameter is twice that of $\Omega(x)$. Some examples are shown in the top-right of Figure 3.
Additionally, as shown in the bottom two rows of Figure 3, we can unfold any portion of the output field into the overlapping regions it contains. The figure shows unfolded regions $\{\Omega_{k}(x)\}$ within two windows of the output field. For each unfolded set we visualize: the windowing functions $w_{k}(x;\mathbf{p})$; the regional boundaries $b_{k}(x;\mathbf{g})$ obtained by applying the nonlinearity in Equation 7 to the local distance functions $d_{k}(x;\mathbf{g})$; and the supports $s_{kj}(x;\mathbf{g})$ colored according to the wedge features $\mathbf{f}_{kj}$ that they gather from the input image. In the top example, the neighborhood is homogeneous, so the windowing functions (via $\mathbf{p}[n]$) have large supports, and there are few boundaries (via $\mathbf{g}[n]$). In the bottom example there is fine-scale geometry. The windowing functions are narrow and the primitives agree on a Y-junction.
4 Network Architecture
We design our network to iteratively refine the fields $\{(\mathbf{g}^{t}[n],\mathbf{p}^{t}[n])\}$. We do so by embedding each field element using a higher dimensional representation, $\boldsymbol{\gamma}^{t}[n]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[n]\in\mathbb{R}^{D_{\pi}}$ respectively, which can be updated via dot-product attention. In practice, we use $D_{\gamma}=64$ and $D_{\pi}=8$, which provides the network with enough capacity to learn meaningful hidden states. We learn simple linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$ which are used for the gather and slice operations. Importantly, all of our network’s elements are invariant to discrete spatial shifts of the image, operating either on individual locations $n$ or on small neighborhoods of locations. See Figure 4 and additional details in the supplement.
Given an input image, the network first applies a neighborhood MLP-mixer, which is a modified variation of MLP-Mixer [32] where we replace the linear operations with convolutions of kernel size $3$. The other change is that we map the input pixels to the hidden state size with a pixel-wise linear mapping rather than taking patches of the input. This block, which we denote the “neighborhood MLP-mixer” transforms the input into an initial hidden state, which is then processed through eight iterations of our boundary attention to further refine the hidden state.
The eight iterations of refinement are broken into two Boundary Attention Blocks with independent weights. To process our input, we first add in a linear projection of the initial hidden state. This is essentially a skip connection that allows our network to retain information from the input pixels at later stages of processing. Next, we copy our hidden state into two identical pieces. We concatenate a dimension 8 learned windowing token to one of the copies and the current estimate of the smoothed global features to the other. We then do neighborhood cross-attention: each pixel in the first copy does two iterations of cross attention with a size 11 patch of the second copy. We add a learned $11\times 1$ 11 positional encoding to the patch, which allows our network to access relative positioning, even if global position cues are absent. We follow each self attention layer with a small MLP.
To transform our output or intermediary hidden state into junction space and render output images, we use a simple linear mapping. We separate the windowing token (the last 8 dimensions) from the hidden state (the first 64 dimensions) and project each through a linear layer. We map the hidden states to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$. These serve as the inputs to our gather and slice operators.
Overall, our network has $2.07\cdot 10^{5}$ learnable parameters, making it orders of magnitude smaller than most learned boundary detectors. For contrast, EDTER [28] has 109 million parameters for its full model.
4.1 Training
We train our network in two stages: we begin by training the neighborhood MLP-mixer with the first Boundary Attention Block, and then we add the second block and retrain end-to-end. We apply our loss to the two final iterations of our network, so for the first stage of training we apply a loss to iterations 3 and 4, and for end-to-end refinement we apply a loss to iterations 7 and 8. We weigh the final loss three times higher than the second to last loss, which encourages the network to allocate capacity to producing high quality outputs, while ensuring that gradient information is shared across the network.
We train our method using a combination of four global losses applied to global (i.e. averaged) fields, and two patch-wise losses applied to individual patches. The first two losses are supervision losses penalizing mismatches between our network’s predictions and the ground truth feature and boundary maps:
$\mathcal{L}_{f}$ $=\sum_{n}\alpha[n]\|\bar{\mathbf{f}}[n]-\mathbf{f}_{\text{GT}}[n]\|^{2},$ (9)
$\mathcal{L}_{d}$ $=\sum_{n}\alpha[n]\left(\bar{d}[n]-d_{\text{GT}}[n]\right)^{2},$ (10)
where $\mathbf{f}_{\text{GT}}$ and $d_{\text{GT}}$ are the ground truth features and distance maps, respectively, and $\alpha[n]$ is a pixel importance function defined as:
$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+C,$ (11)
with $\beta$ and $C$ controlling how much weight to give pixels near boundaries. We set $\beta=0.1$, $\delta=1$, and increase $C$ throughout training to give more weight to non-boundary locations. We also tested a more involved pixel importance mask that gave extra weight to visible vertices and intersections, the details for which can be found in the supplement. Note that using noiseless feature maps for supervision in Equation 9 encourages the windowing functions to be large in smooth regions.
On top of the two supervision losses we apply two consistency losses from [33], that minimize the per-pixel variances $\nu_{f}[n]$ and $\nu_{d}[n]$. Similar to the supervision losses, we weigh those by $\alpha$ from Equation 11. These consistency losses encourage the junction shapes $\mathbf{g}$ in overlapping regions to agree. Minimizing $\nu_{f}[n]$ also encourages windowing functions to be large, because that increases the gather area, which in turn reduces noise in wedge features $\mathbf{f}_{nj}$ that are sliced to compute the variance $\nu_{f}[n]$ at each $n$.
Finally, we use two patch-wise losses to encourage individual feature and distance patches to agree with the supervisory ones:
$\ell_{f}=\sum_{k}\chi[k]\sum_{n\in\Omega_{k}}\alpha[n]\|\bar{\mathbf{f}}[n]-\mathbf{f}_{\text{GT}}[n]\|^{2},$ (12)
$\ell_{d}=\sum_{k}\chi[k]\sum_{n\in\Omega_{k}}\alpha[n](\bar{d}[n]-d_{\text{GT}}[n])^{2},$ (13)
where $\chi[k]$ is a patch importance function defined as:
$\chi[k]=\left(\sum_{n\in\Omega_{k}}(d_{\text{GT}}[n]+\delta^{\prime})\right)^{-1},$ (14)
and we set $\delta^{\prime}=1$. These per-patch losses gives the network a more direct signal on how to adjust its weights than purely global losses, which average over multiple patches.
Figure 5: The ODS F-score of our method at different noise levels, compared with the Field of Junctions (FOJ) [33] run using patch sizes $9\times 9$, $17\times 17$, and $25\times 25$, and EDTER [28], with and without preprocessing by BM3D [9]. The bottom insets show example patches at representative PSNR values.
5 Experiments
Implementation details.
We train our model on noisy synthetic data of colorful combinations of triangles and circles. We render $240\times 320$ images containing $15$ to $20$ shapes each, but use $125\times 125$ crops for training. To those crops we add Gaussian and/or Perlin noise [26], and with probability $0.1$ we average over the color channels to produce grayscale inputs. Our dataset contains $10^{5}$ images, $90\%$ of which are used for training, and the rest for testing. For training and optimization details, refer to the supplement.
Performance w.r.t. noise levels.
Figure 5 shows the comparison of our method and baseline approaches, the Field of Junctions [33] and EDTER [28] under different noise levels. We test the Field of Junctions with different patch sizes ($9\times 9$, $17\times 17$, and $25\times 25$) and EDTER with/without preprocessing using optimally-tuned denoising using BM3D [9]. The tuneable parameters for Field of Junctions were chosen to maximize its performance on noisy images with $17\times 17$ patches. Our method outperforms all baselines at lower noise levels and is competitive with the Field of Junctions at higher noise levels, while being orders of magnitude faster than it (see Table 1).
Figure 6: Our network can reliably locate boundaries with sub-pixel precision, due in part to our parametric representation for boundaries. We calculate the ODS F-Score of synthetic images as we increase the pixel matching distance threshold (how close a pixel must be to a ground truth pixel to be counted) from 0 to 1.5 and vary the PSNR of the images from $\infty$ to 8.0. Each line represents the performance at a noise level, with lighter-colored lines corresponding to noisier inputs. X-axis values below 1.0, denoted by the vertical dotted line, represent sub-pixel precision.
Figure 7: One of the advantages of our parametric form for boundaries is that we can trivially upsample our output boundary maps by striding the predicted junction parameters, and scaling the patch size accordingly. This results in clean boundaries even when our output boundary images are upsampled from $125\times 125$ (left) to $500\times 500$ (right).
Figure 8: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$. Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$. The embedding has learned to be smooth and have an intuitive zero.
Sub-pixel precision.
To measure sub-pixel precision, we render $500\times 500$ images containing pairs of overlapping circles and triangles. This gives us a precise binary boundary map for that resolution. We downsample the images to $125\times 125$, and add varying amounts of Gaussian noise. These noisy, downsampled images serve as the inputs. To evaluate the predicted boundaries, we upsample the outputs back to $500\times 500$ pixels. For EDTER, we use bilinear interpolation along with edge thinning to produce the upsampled boundaries. In our case, a byproduct of our parametric form for boundaries is that an intuitive method for upsampling boundaries naturally arises: by increasing the patch stride to $4$, and rendering the patches as $4$ times as large—68×$\times 68$ 68 pixels—we can create an upsampled version of our image while retaining our method’s accuracy in boundary localization. We visualize this process in Figure 7.
We evaluate the upsampled outputs on the original $500\times 500$ binary ground truth maps and vary the minimum distance threshold of the evaluation metric, which denotes how close a prediction must be to the ground truth to be used. The results are shown in Figure 6. Our F-score remains high across all matching thresholds, even when the inputs contain additive Gaussian noise.
Figure 9: Evolution of boundaries during iterations, in reading order. Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
Figure 10: Our model generalizes well to real images, and can find accurate boundaries in natural images. Top left: The input image, a low light image from the SIDD [1] dataset. Top right: Our network’s predicted boundaries. Bottom left: The output mean feature map. Bottom right: The predicted distance map.
Linear interpolation in junction space.
Surprisingly, we note that our network learns a spatially smooth manifold of junctions in its hidden state. Figure 8 visualizes some properties of its learned embedding of junction space. We generate equally-spaced samples $\boldsymbol{\gamma}_{i}\in\mathbb{R}^{D_{\boldsymbol{\gamma}}}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to 00 and then to a particular $\boldsymbol{\gamma}_{b}$. We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center. For visual comparison, we show an analogous interpolation from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$ by using a linear geometric interpolation in junction space ${\cal G}$, the expressions for which are in the supplement.
Evolution of the outputs over time.
Figure 9 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement. Specifically, we visualize the result of slicing similar to Equation 5 but with the regional distance functions $d_{k}$ replaced by their nonlinear counterparts $b_{k}$. We see that early iterations are exploratory and unstructured, and that later iterations reach agreement.
Results on real images.
Figures LABEL:fig:real_results and 10 show the qualitative results on real images. Despite being trained on synthetic data, our method can outperform existing state-of-the-art methods at multiple levels of real sensor noise present in ELD [34]. Our method produces crisp and well-defined boundaries despite high levels of noise.
Table 1: Inference time (in seconds) of EDTER [28] (without preprocessing by BM3D [9]), and Field of Junctions (FoJ) [33] run at various patch sizes, compared with our method. We compare both with the original (FoJ) implementation as well as our JAX reimplementation (JAX-FoJ). Unlike FoJ [33], our method does not require striding for fitting moderately-sized images in memory (runtimes reported with $*{}^{*}$ did require striding). The average run times were calculated on an Nvidia A100 GPU.
$125\times 125$ $320\times 320$
EDTER [28] $0.130$ $0.130$
FoJ [33]  $9\!\!\times\!\!9$ $36.8$ $206$
 $17\!\!\times\!\!17$ $88.0$ $76.0^{*}$ (stride-$3$)
 $25\!\!\times\!\!25$ $162$ $57.6^{*}$ (stride-$5$)
[…]
Ours $0.0823$ $0.678$
Inference time.
Table 1 compares the running time for different methods at two resolutions. We denote several runs for Field of Junctions where we had to increase the patch stride (optimize junctions at strided intervals) to avoid running out of memory. Notably, we are orders of magnitude faster than Field of Junctions, despite achieving similar performance. Additionally, because our network can handle inputs of any size, the run time depends on the input shape, in contrast with EDTER that has a maximal input size of $320\times 320$ inputs and constant run time.
6 Conclusion
We have introduced a differentiable model that explicitly reasons about geometric primitives such as edges, corners, junctions, and regions of uniform appearance, by using boundary attention. Despite being trained on simple geometric synthetic data, our method generalizes to natural images, and predicts clean boundaries, even when the images it is applied to are significantly corrupted by noise. Furthermore, the parametric form of our model elicits a natural formulation for boundaries with subpixel precision, and its bottom-up patch-wise approach enables it to process any image at its native resolution and aspect ratio.
Supplemental Material
Boundary Attention: Learning to Find Faint Boundaries at Any Resolution
1 . The space of $M$-junctions
Here we provide the expressions for the support functions $s_{j}(x;\mathbf{g})$ and the unsigned distance function $d(x;\mathbf{g})$ from Section 3 of the main paper. We also describe the differences between our parameterization of junction space and the original one in the field of junctions [33], with the new parameterization’s main advantages being the avoidance of singularities and the ability to define mechanisms for smooth interpolation. Our descriptions of these require introducing a few additional mathematical details. We provide these details for the general case of geometric primitives (junctions) $\mathbf{g}$ that have $M$ angular wedges $\boldsymbol{\omega}=(\omega_{1},\ldots,\omega_{M})$, for which the paper’s use of $M=3$ is a special case.
To begin, consider a local region $\Omega(x)\subset\mathbb{R}^{2}$ and fix a positive integer value for the maximum number of angular wedges $M>0$ (the paper uses $M=3$). Our partitions are parameterized by $\mathbf{g}\in\mathbb{R}^{2}\times\mathbb{S}^{1}\times\triangle^{M-1}$, where $\mathbb{S}^{1}$ is the unit circle and $\triangle^{M-1}$ is the standard $(M-1)$-simplex (i.e., the set of $M$-vectors whose elements are nonnegative and sum to one). We use the notation $\mathbf{g}=(\boldsymbol{u},\theta,\boldsymbol{\omega})$, where $\boldsymbol{u}=(u,v)\in\mathbb{R}^{2}$ is the vertex, $\theta\in\mathbb{S}^{1}$ is the orientation, and $\boldsymbol{\omega}=(\omega_{1},\omega_{2},\ldots,\omega_{M})$ are barycentric coordinates (defined up to scale) for the $M$ relative angles, ordered clockwise starting from $\theta$. As noted in the main paper, our convention is to express the vertex coordinates relative to the center of region $\Omega(x)$, and we note again that the vertex is free to move outside of this region. We also note that up to $M-1$ of the angles $\omega_{j}$ can be zero. When necessary, we use notation $\tilde{\boldsymbol{\omega}}=(\tilde{\omega}_{1},\tilde{\omega}_{2},\ldots,\tilde{\omega}_{M})$ to represent angles that are normalized for summing to $2\pi$:
$\tilde{\boldsymbol{\omega}}=\frac{2\pi\boldsymbol{\omega}}{\sum_{j=1}^{M}\omega_{j}}.$ (15)
As an aside, we note that there are some equivalences in this parameterization. First, one can perform, for any $k\in\{1\ldots(M-1)\}$, a cyclic permutation of the angles $\boldsymbol{\omega}$ and adjust the orientation $\theta$ without changing the partition. That is, the partition does not change under the cyclic parameter map
$\omega_{j}$ $\rightarrow\omega_{j+k(\textrm{mod}M)}$ (16)
$\theta$ $\rightarrow\theta-\sum_{j=M+1-k}^{M}\omega_{j}$ (17)
for any $k\in\{1\ldots(M-1)\}$. Also, an $M$-junction $\left(\boldsymbol{u},\theta,(\omega_{1},\ldots,\omega_{M})\right)$ provides the same partition as any $M^{\prime}$-junction, $M^{\prime}>M$, that has the same vertex and orientation along with angles $(\omega_{1}\ldots\omega_{M},0\dots)$. This captures the fact that $M$-junction families are nested for increasing $M$.
Figure 11: Anatomy of an $M$-junction $\mathbf{g}=(\mathbf{u},\theta,\boldsymbol{\omega})$ with $M=3$. Left: Boundary directions $\phi_{j}$ and central directions $\psi_{j}$ are determined directly from relative angles $\boldsymbol{\omega}$ and orientation $\theta$ (which is equal to $\phi_{1}$). Middle panels: Unsigned distance function for a boundary ray $d_{3}(x;\mathbf{g})$ and overall unsigned distance function $d(x;\mathbf{g})$, which is the minimum of the three per-ray ones. Right: Associated boundary function $b_{\eta}(x;\mathbf{g})$ using $\eta=0.7$.
As shown in Figure 11, other geometric features of a junction can be directly derived from the orientation and angles. The central directions $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{M})$ are
$\psi_{j}=\theta+\frac{\tilde{\omega}_{j}}{2}+\sum_{k=1}^{j-1}\tilde{\omega}_{k},\quad j\in\{1\ldots M\},$ (18)
and the boundary directions $\boldsymbol{\phi}=(\phi_{1},\ldots,\phi_{M})$ are given by $\phi_{1}=\theta$ and
$\phi_{j}=\theta+\sum_{k=1}^{j-1}\tilde{\omega}_{k},\quad j\in\{2\ldots M\}.$ (19)
A key difference between our new parameterization of $M$-junctions and the original one [33] is that the latter comprises $\left(\boldsymbol{u},\boldsymbol{\phi}\right)$ and requires enforcing constraints $0\leq\phi_{1}\leq\phi_{2}\leq\cdots\leq\phi_{M}\leq 2\pi$ (or somehow keeping track of the permutations of wedge indices that occur when these constraints are not enforced). The new $\left(\boldsymbol{u},\theta,\boldsymbol{\omega}\right)$-parameterization eliminates the need for such constraints.
As noted in the main paper’s Section 3, we define the $j$th support $s_{j}(x;\mathbf{g})$ as the binary-valued function that indicates whether each point $x\in\Omega$ is contained within wedge $j\in\{1\ldots,M\}$. Its expression derives from the inclusion condition that the dot product between the vector from the vertex to $x$ and the $j$th central vector $\left(\cos\psi_{j},\sin\psi_{j}\right)$ must be smaller than the cosine of half the angle $\tilde{\omega}_{j}$. Using Heaviside function $H(\cdot)$ we write
$\begin{split}s_{j}(x;\mathbf{g})=H\Big{(}(x-\mathbf{u})\cdot(\cos\psi_{j},\sin\psi_{j})-\cos(\tilde{\omega}_{j}/2)||x-\mathbf{u}||\Big{)}.\end{split}$ (20)
As an aside, observe that this expression remains consistent for the case $M=1$, where there is a single wedge. In this case, $\tilde{\boldsymbol{\omega}}=\tilde{\omega}_{1}=2\pi$ by Equation 15, and the support reduces to $s_{1}(x)=1$ for all vertex and orientation values.
The unsigned distance $d(x;\mathbf{g})$ represents the Euclidean distance from point $x$ to the nearest point in the boundary set defined by $\mathbf{g}$. It is the minimum over $M$ sub-functions, with each sub-function being the unsigned distance from a boundary ray that extends from point $\mathbf{u}$ in direction $\phi_{j}$. The unsigned distance from the $j$th boundary ray is equal to the distance from its associated line for all points $x$ in its containing half-plane; and for other points it is equal to the radial distance from the vertex. That is,
$d_{j}(x;\mathbf{g})=\begin{cases}\left|(x-\mathbf{u})\cdot(-\sin\phi_{j},\cos\phi_{j})\right|,&\text{if }(x-\mathbf{u})\cdot(\cos\phi_{j},\sin\phi_{j})>0\\ \|(x-\mathbf{u})\|,&\text{otherwise}.\end{cases}$ (21)
Then, the overall distance function is
$d(x;\mathbf{g})=\min_{j\in{1\dots M}}d_{j}(x;\mathbf{g}).$ (22)
Finally, analogous to Equation 7 in the main paper, we define a junction’s boundary function $b_{\eta}(x;\mathbf{g})$ as the result of applying a univariate nonlinearity to the unsigned distance:
$b_{\eta}(x;\mathbf{g})=\left(1+(d(x;\mathbf{g})/\eta)^{2}\right)^{-1}.$ (23)
Figure 11 shows an example of a junction’s distance function and its associated boundary function with $\eta=0.7$.
Interpolation
Another advantage of the present parameterization compared to that of the original [33] is that it is a simply-connected topological space and so allows for defining mechanisms for smoothly interpolating between any two junctions $\mathbf{g}$ and $\mathbf{g}^{\prime}$. In our implementation we simply define interpolation variable $t\in[0,1]$ and compute interpolated junctions $\mathbf{g}^{(t)}=\{\boldsymbol{u}^{(t)},\theta^{(t)},\boldsymbol{\omega}^{(t)}\}$ using
$\boldsymbol{u}^{(t)}$ $=(1-t)\boldsymbol{u}+t\boldsymbol{u}^{\prime}$ (24)
$\tilde{\boldsymbol{\omega}}^{(t)}$ $=(1-t)\tilde{\boldsymbol{\omega}}+t\tilde{\boldsymbol{\omega}}^{\prime},$ (25)
and
[…]
$(p,q)$ $=\operatorname{Slerp}\left((\cos\theta,\sin\theta),(\cos\theta^{\prime},\sin\theta^{\prime}),t\right),$
where $\operatorname{Slerp}()$ is the 2D geometric spherical linear interpolation operator,
$\operatorname{Slerp}(\boldsymbol{p},\boldsymbol{p}^{\prime},t)=\frac{\sin\left((1-t)\Delta\theta\right)}{\sin\left(\Delta\theta\right)}\boldsymbol{p}+\frac{\sin\left(t\Delta\theta\right)}{\sin\left(\Delta\theta\right)}\boldsymbol{p}^{\prime},$
with $\Delta\theta=\arccos(\boldsymbol{p}\cdot\boldsymbol{p}^{\prime})$. The bottom row of Figure 9 in the main paper visualizes a set of samples from smooth trajectories in junction space using this mechanism.
2 . Training Data
Figure 12: Columns 1 to 5: Examples of the synthetic data used to train our model using supervision with ground-truth boundaries. Column 6: Rendered distance maps corresponding to column 5. The training data contains random circles and triangles that each have a random RGB color, and the images are corrupted by various types and amounts of noise. Each noiseless image has an unrasterized, vector-graphics representation of its shapes and colors, which specify the clean image and exact boundary-distance map with unlimited resolution.
We find that we can train our model to a useful state using purely synthetic data, examples of which are depicted in Figure 12. In fact, we find it sufficient to use very simple synthetic data that consists of only two basic shapes—circles and triangles—because these can already produce a diverse set of local edges, thin bars, curves, corners, and junctions, in addition to uniform regions. We generate an image by randomly sampling a set of circles and triangles with geometric parameters expressed in continuous, normalized image coordinates $[0,1]\times[0,1]$. We then choose a random depth ordering of the shapes, and we choose a random RGB color for each shape. Importantly, the shape and color elements are specified using a vector-graphics representation, and the shape elements are simple enough to provide an exact, symbolic expression for each image’s true boundary-distance map, without approximation or rasterization. They also allow calculating the precise locations, up to machine precision, for all of the visible corners and junctions in each image.
At training time, an input image is rasterized and then corrupted by a random amount and type of noise, including some types of noise that are spatially-correlated. This forces our model to only use color as its local cues for boundaries and grouping; and it forces it to rely heavily on the topological and geometric structure of curves, corners and junctions, as well as their contrast polarities. The highly-varying types and amounts of noise also encourages the model to use large window functions $w(x;\mathbf{g})$ when possible, since that reduces noise in the gather operation and reduces variance $\nu_{f}[n]$.
Our dataset, which we call Kaleidoshapes, will be released upon publication, along with the code for generation, training and evaluation.
Shapes and colors. For our experiments, we rasterized each image and its true distance map at a resolution of $240\times 320$ images, with each one containing between 15 and 20 shapes. We used a $40\!\!:\!\!60$ ratio of circles to triangles. In terms of normalized coordinates, circles had radii in the range $[0.05,0.2]$ and triangles had bases in the range $[0.02,0.5]$ and heights in the range $[0.05,0.3]$. This allows triangles to be quite thin, so that some of the local regions $\Omega(x)$ contain thin bar-like structures. Additionally, we included a minimum visibility threshold, filtering out any shapes whose visible number of rasterized pixels is below a threshold. Colors were selected by uniformly sampling all valid RGB colors. During training, batches consisted of random $125\times 125$ crops.
Noise. For noise types, we used combinations of additive zero-mean Gaussian noise; spatially average-pooled Gaussian noise; Perlin noise [26], and simulated photographic sensor noise using the simplified model from [34]. The total noise added to each image was sampled uniformly to be between 30% and 80% of the maximum pixel magnitude, and then noise-types were randomly combined with associated levels so that they produced the total noise level. Since zero-mean noise can at times result in values below 0 or above the maximum magnitude threshold, we truncate any pixels outside of that range.
3 . Model Details
Our model is designed to be purely local and bottom up, with all of its compositional elements operating on spatial neighborhoods in a manner that is invariant to discrete spatial shifts of an image. Its design also prioritizes having a small number of learnable parameters. Here we provide the details of the two blue blocks in the main paper’s Figure 3: Neighborhood MLP-Mixer and Neighborhood Cross-attention. We implement our model in JAX, and we will publicly share our model code and its pretrained weights upon publication.
S3.1 . Neighborhood MLP-Mixer
Our neighborhood MLP-mixer is a shift invariant, patch-based network inspired by MLP-mixer [32]. It replaces the image-wide operations of [32] with patch-wise ones. Given an input image, we first linearly project its pixels from $\mathbb{R}^{3}$ to dimension $\mathbb{R}^{D_{\gamma}}$ (we use $D_{\gamma}=64$), which is followed by two neighborhood mixing blocks. Each neighborhood mixing block contains a spatial patch mixer followed by a channel mixer. The spatial patch mixer is implemented as two $3\times 3$ spatial convolutions with weights tied across channels. It thereby combines spatial patches of features with all channels (and patches) sharing the same weights. Following [32], we use GELU [17] activations. The channel mixer is a per-pixel MLP with spatially-tied weights. To handle border effects in our neighborhood MLP-mixer, we apply zero-padding after the initial projection from $\mathbb{R}^{3}$ to $\mathbb{R}^{64}$, and then we crop to the input image size after the second neighborhood mixing block to remove features that correspond to patches without full coverage, i.e., patches that contain pixels outside of the original image.
S3.2 . Neighborhood Cross-attention
The neighborhood cross-attention block similarly enforces shift-invariance and weight sharing across spatial neighborhoods. Inside this block are two transformer layers whose cross-attention components are replaced with neighborhood cross-attention components that are restricted to a spatial neighborhood of pixels. We use $11\times 1$ 11 neighborhoods in our implementation. In each neighborhood containing a query token, we add a learned positional encoding to the key/value tokens which is relative to the neighborhood’s center and is the same for all neighborhoods. Then the query is updated using standard cross-attention with its neighborhood of key/values. We use 4 cross-attention heads. Like the standard transformer, each neighborhood cross attention component is followed by an MLP, dropout layer, and additive residual. To handle border effects, we zero-pad the key and value tokens so that every query attends to an $11\times 1$ 11 neighborhood, and then zero-out any attention weights involving zero-padded tokens.
S3.3 . Training details
We pretrain the neighborhood MLP-mixer and the first boundary attention block on a simplified variation of our Kaleidoshapes dataset, where each image is $100\times 100$ and contains a single triangle and circle with additive zero-mean Gaussian noise. We omit the global losses of Equations $9$ and $10$ during this pretraining phase. This primes the network to learn meaningful hidden states $\boldsymbol{\gamma}[n]$ and prevents the “collapsing” of junctions, where the boundary-consistency loss (i.e. the sum over pixels of variance of distance $\nu_{d}[n]$) dominates and the network learns to predict all-boundaryless patches that are globally consistent but inaccurate. Because of data imbalance—only a small fraction of regions $\Omega_{n}(x)$ contain corners or junctions—we add an additional spatial importance mask to prioritize the regions that contain a corner (i.e., a visible triangle vertex) or a junction (i.e., an intersection between a circle and a triangle’s edge). Our data generation process produces a list of all non-occluded vertices and intersections in each image, and we use these values to create a spatial importance mask with gaussians centered at each of these points. In practice, we use gaussians with a standard deviation of 7 pixels. This mask is added to the loss constant $C$.
The final stage of training adds a second boundary attention block with weights that are initialized using a copy of the pretrained weights of the first boundary attention block. We use $100,000$ crops of size $125\times 125$ from our Kaleidoshape images (10% withheld for testing) and the full set of losses; and we optimize all of the model’s parameters, including those of the neighborhood MLP-mixer and the first boundary attention block. Like in pretraining, we add a spatial importance that prioritizes region containing a corner (i.e., a visible triangle vertex) or a junction (i.e., a visible intersection between the boundaries of any two shapes).
Our trained weights will be released upon publication.
4 . Qualitative Behavior for Natural Images
In Figures 14 and 13, we show how the model behaves on noiseless natural images that contain texture and recognizable objects. In particular, Figure 13 emphasizes how the boundary maps produced by our model qualitatively differ from those of many classical bottom-up edge-detectors and also from those of learned, end-to-end models that have been trained to match human annotations. The figure compares our output to that from Canny [4], Field of Junctions [33] with a patch size of $11$, Pidinet [31], and EDTER [28]. The latter two methods are trained on human annotated data, whereas the former two methods, like our model, are not. (Note that inputs for all models besides EDTER [28] were $300\times 400$. Input to EDTER was down-sampled to $225\times 300$ due to its input size constraint.)
We find that our model produces finer structures than the end-to-end learned models [10, 28] because it is trained to only use local spatial averages of color as its cue for boundaries and grouping. It does not include mechanisms for grouping based on local texture statistics, nor based on non-local shape and appearance patterns that have semantic meaning to humans. Compared to the bottom-up methods of Canny [4] and Field of Junctions [33], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of field $\mathbf{p}[n]$. In contrast, the Field of Junctions and Canny both operate at a single pre-determined choice of local size, so they tend to oversegment some places while undersegmenting others.
Figure 13: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [28] and Pidinet [31] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [4], and Field of Junctions [33] with patch size $11$.
Figure 14: Qualitative behavior of our model on noiseless natural images. From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
5 . Additional Examples for Low-light Images
Figure 15 shows examples of applying our model to indoor images taken by an iPhone XS in low light conditions.
Figure 15: Visualization of our model’s output for low-light images captured by an iPhone XS. From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
Figure 16 provides additional comparisons for a sample of varying-noise images from the ELD dataset [34]. When detecting boundaries at low signal-to-noise ratios, it is difficult to accurately discern finer structures as the noise level increases. Some algorithms, such as Field of Junctions [33], have tunable parameters such as patch-size that provide control over the level of detection. A small patchsize allows recovering fine structures in lower noise situations, but it causes many false positive boundaries at high noise levels. Conversely, a large patchsize provides more resilience to noise but has not ability to recover fine structure at all. Our model reduces the severity of this trade-off by automatically adapting its local windowing functions in ways that have learned to account for both the amount of noise and the local geometry of the underlying boundaries.
In Figure 16 we see that our model is able to capture the double-contour shape of the curved, thin black bars, and that it continues to resolve them as the noise level increases, more than the other low-level methods. We also note that only the low-level models resolve this level of detail in the first place: The models trained on human annotations—EDTER, HED, Pidinet, and Structured Forests—miss the double contour entirely, estimating instead a single thick curve. We emphasize again that a user can adjust the behavior of Canny and Field of Junctions by tuning their local size parameters, either the filter size for Canny or the patchsize for Field of Junctions. Increasing the local size improves their resilience to noise but reduces their spatial precision. Neither system provides the ability to estimate fine grained details and withstand noise, like our model does.
Figure 16: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise. We compare to end-to-end models that are trained to match human annotations (EDTER [28], HED [35], Pidinet [31], and Structured Forests [10]) in addition to low-level models that are not (Canny [4], and Field of Junctions (FOJ) [33]).
Figure 20 contains additional examples of images cropped from the ELD dataset. Here we include examples with even higher levels of noise to show the complete degradation of our algorithm and others.
Figure 17:
Figure 18: (cont.)
Figure 19: (cont.)
Figure 20: (cont.) Additional qualitative comparisons between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, using crops from the ELD dataset under increasing amounts of photographic noise, including very high levels of noise.
6 . Additional Uses of Our Model
Here we demonstrate to uses of our model that follow directly from its output: hole-filling in RGBD images and non-photorealistic stylization.
S6.1 . Color-based Depth Completion
Figure 21 shows an example of using our model for simple hole-filling in the depth channels of RGBD images from the Middlebury Stereo Datasets [30, 29]. We run our model on the RGB channels, and then for each pixel $n$ that has a missing depth value, we use our model’s output local attention kernels $a_{n}(x)$ to fill in that pixel’s value using an attention-weighted average of the observed depth values around it. This simple algorithm can be applied whenever the hole sizes are smaller than the maximum diameter of our attention maps, which is $34\times 34$ in our current implementation).
Figure 21: Using our model for depth completion in RGBD images. Left: Input RGB channels. Middle: Input depth channel, with dark blue indicating missing values. Right: Completed depth using our model’s output attention kernels.
S6.2 . Application: Photo Stylization
Figure 22 shows examples of using our model’s output for image stylization, by superimposing an inverted copy of the output boundary map $\bar{b}_{\eta}[n]$ onto the smoothed colors $\bar{\mathbf{f}}[n]$.
Figure 22: Examples of stylized natural photographs, created by imposing our method’s output boundary map onto the output smoothed colors.

### v2

@@
-Boundary Attention: Learning to Find Faint Boundaries at Any Resolution
+Boundary Attention: Learning to Localize Boundaries under High Noise
 Abstract
-We present a differentiable model that explicitly models boundaries—including contours, corners and junctions—using a new mechanism that we call boundary attention.
-We show that our model provides accurate results even when the boundary signal is very weak or is swamped by noise.
-Compared to previous classical methods for finding faint boundaries, our model has the advantages of being differentiable; being scalable to larger images; and automatically adapting to an appropriate level of geometric detail in each part of an image.
-Compared to previous deep methods for finding boundaries via end-to-end training, it has the advantages of providing sub-pixel precision, being more resilient to noise, and being able to process any image at its native resolution and aspect ratio.
+We present a differentiable model that infers explicit boundaries, including curves, corners and junctions, using a mechanism that we call boundary attention.
+Boundary attention is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specify an unrasterized description of the local boundary structure in every overlapping patch within an image.
+It operates in a bottom-up fashion, similar to classical methods for sub-pixel edge localization and edge-linking, but with a higher-dimensional description of local boundary structure, a notion of spatial consistency that is learned instead of designed, and a sequence of operations that is end-to-end differentiable.
+We train our model using simple synthetic data and then evaluate it using photographs that were captured under low-light conditions with variable amounts of noise.
+We find that our method generalizes to natural images corrupted by real sensor noise, and predicts consistent boundaries under increasingly noisy conditions where other state-of-the-art methods fail.
 1 Introduction
-Converting a precise contour that is defined mathematically in continuous 2D space to its discrete representation in pixel-space is a common task in computer graphics, and there are established techniques for rasterization [12, 27], anti-aliasing [16, 6] and so on.
-However, the inverse problem in computer vision of robustly inferring precise and unrasterized descriptions of contours from discrete images remains an open challenge.
-Earlier work in the field inferred unrasterized representations of edges [4, 8, 25, 14, 13, 22]—and sometimes corners and junctions too—and it explored algorithms for connecting them into boundaries via edge-linking and MRFs.
-But since the dawn of deep learning, boundaries have almost exclusively been modeled using discrete, rasterized maps.
-Moreover, it is not uncommon to rely heavily on supervision from human annotations, and even modern datasets can suffer from annotation errors on the order of pixels, quite far from sub-pixel accuracy.
-Taking inspiration from early computer vision work, we present a model for inferring unrasterized boundaries that can be learned instead of designed.
-We show that this model is able to benefit from the power of deep learning while also realizing many of the advantages—such as robustness to noise, sub-pixel precision, and adaptability between signal-types—that have long been the potential strengths of classical bottom-up techniques.
+Converting a precise contour that is defined mathematically in continuous 2D space to a discrete pixel representation is a common task in computer graphics, and there are established tools for rasterization [8, 18], anti-aliasing [11, 4] and so on.
+However, the inverse problem in computer vision of robustly inferring precise, unrasterized contours from discrete images remains an open challenge, especially in the presence of noise.
+Earlier work by Canny [3] and many others [5, 16, 10, 9, 15] explored the detection of parametric edge models, and there are a variety of bottom-up algorithms that try to connect them by encouraging geometric consistency via edge-linking or message-passing.
+But since the dawn of deep learning, boundaries have almost exclusively been represented using discrete, rasterized maps, and spatial-consistency mechanisms that were previously based on explicit curve geometry have largely been replaced by black-box neural networks.
+In this paper, we take inspiration from early computer vision work and revisit the task of finding unrasterized boundaries via bottom-up geometric processing.
+But unlike early work, we want to leverage the speed and power of neural networks to learn these processes instead of hand-crafting them.
+We focus our attention on the low-level task of finding boundaries that separate regions of uniform color, as depicted by the toy problem in Figure 1.
+This task becomes difficult at high noise levels, especially within local patches, and we address it by creating a model that can learn to exploit a wide range of low-level cues, such as curves being predominantly smooth with low-curvature; curves tending to meet at corner and junction points; contours tending to have consistent contrast polarity throughout their extent; colors tending to vary smoothly at locations away from boundaries; and so on.
 The core of our model is a mechanism we call boundary attention.
-It is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specifies the local boundaries surrounding every pixel.
-The model’s output is a field of overlapping geometric primitives that can be used in a variety of ways, including to produce an unsigned distance function for the image’s boundaries, a boundary-aware smoothing of its channel-values, and a field of soft local attention maps that relate every pixel to its neighbors.
-To evaluate our model, we consider the problem of finding boundaries in images corrupted by extreme amounts of noise [23, 24, 33], where the images can be of any size and resolution, and where we know little about the noise and the objects or shapes we are looking for.
-We choose this problem because it relies entirely on having a strong model for the basic topological and geometric properties of boundaries, namely that they are piecewise smooth curves which connect at corners or junctions and divide regions of homogeneity.
-Figure 1: Our model uses neighborhood attention with dense, stride-1 tokens (but we depict non-overlapping ones here for visual clarity).
-The entire model is invariant to discrete spatial shifts and so applies at any resolution.
-Each token encodes an adaptively-sized geometric primitive that represents the unrasterized local boundaries surrounding a pixel, and the tokens evolve through boundary attention to be geometrically consistent.
-The output is an overlapping field of primitives that directly implies a boundary-aware smoothing of the input image and an unsigned distance map to the image boundaries.
-We show that our model has several advantages.
-All of its components are local and invariant to discrete spatial shifts, so it can be trained on small-sized images and then deployed on much larger and differently-shaped ones.
-It is also very compact, comprising only 207k parameters, and it runs several times faster than many alternatives.
-We also find that it can be trained to a useful state with very simple synthetic data, made up of random circles and triangles that are uniformly colored and then corrupted by noise.
-Despite this simplicity, we find that the model can generalize to real images surprisingly well.
+It is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specify the local boundaries within dense (stride-1) patches of an image.
+The model’s output is a dense field of unrasterized geometric primitives that, as depicted in the right of Figure 1, can be used in a variety of ways, including to produce an unsigned distance function and binarized map for the image boundaries, a boundary-aware smoothing of the input image colors, and a map of spatial affinities between any query point and the pixels that surround it.
+An important feature of our model is that its output patch primitives (see Figure 2) have enough flexibility to represent a wide range of local boundary patterns and scales, including thin bars, edges, corners and junctions, all with various sizes, and all without rasterization and so with unlimited resolution.
+This gives our model enough capacity to learn to localize boundaries with high precision, and to regularize them in the presence of noise without erroneously rounding their corners or missing their junction points.
+Figure 1: Process overview.
+The image unfolds into stride-1 patches, and boundary attention operates iteratively on their embeddings to produce for each patch: (i) a parametric three-way partitioning, and (ii) a parametric windowing function that defines its effective patch size.
+(Figure 2 shows parameterization details.) This output field implies a variety of global maps, shown in clockwise order: a boundary-aware smoothing of the input colors; an unsigned boundary-distance map; a boundary map; and a map of spatial affinities between any query point and its neighbors.
+We design our model so that all of its components are local and invariant to discrete spatial shifts, enabling it to be trained on small-sized images and then deployed on much larger and differently-shaped ones.
+By tying the weights across multiple portions of the network, we increase its resilience to noise which we validate with ablations.
+Our resulting model is very compact, comprising only 207k parameters, and it runs several times faster than many alternatives.
+Further, our model can be trained to a useful state with very simple synthetic data, made up of random circles and triangles that are uniformly colored and then corrupted by noise.
+Despite this simplicity, we find that our model can generalize to captured photographs and novel noise processes surprisingly well.
+We evaluate our model on real noisy images and show that it can capture fine details that other methods miss, and moreover maintain its predictions over high levels of noise.
 Our main contributions can be summarized as follows:
@@
   2. 2.
-    We demonstrate that our model based on boundary attention is more effective than current state-of-the-art methods at finding boundaries at high levels of sensor noise (Fig.
-LABEL:fig:real_results) and can achieve sub-pixel accuracy, even in the presence of noise.
+    We demonstrate that our model based on boundary attention is more effective than current state-of-the-art methods at precisely and consistently finding boundaries at high levels of sensor noise (Figure 6).
+Figure 2: Parameterization details.
+Left: Each patch $k$ is associated with an unrasterized three-way partitioning of its area (colored blue, orange and purple here).
+The partitioning parameters comprise a vertex $(u,v)$, orientation $\theta$, and angles $(\omega_{1},\omega_{2},\omega_{3})$, defined up to scale.
+A: A walk through junction space by linearly interpolating between junctions is spatially smooth, and can represent edges, bars, corners, Y-junctions, T-junctions and uniform regions.
+B: Each junction is modulated through a learned windowing function.
+The windowing parameters $\mathbf{p}=(p_{1},p_{2},p_{3})$ are convex weights over a dictionary of binary pillboxes.
+See the supplement for an interactive visualization.
 2 Related Work
-It is possible to process the outputs of certain filters to localize edges with sub-pixel precision (e.g.,  [4, 8, 25, 14, 13, 22]).
-But this approach struggles near corners and junctions because the edge filters are not derived for these, and accuracy breaks down.
-Recently, the field of junctions [33] showed improvements by expanding the library of local, unrasterized geometric primitives from classical edge-based primitives to a larger family of “generalized junctions”.
-This allows unifying a variety of bottom-up cues that had often previously been handled separately, such as geometrically-consistent and polarity-consistent edge elements being linked into contours [18, 23, 24, 4]; contours that approach one another being joined at corners and junctions [15, 21, 36]; homogeneous regions that look the same being together; and those that look different being apart [7, 11, 2].
-The field of junctions showed that putting all of these cues together leads to a dramatic improved in noise resilience, and it also allows analyzing the recovered boundaries explicitly into its component junctions, edges and corners.
-In this paper, we leverage all of these benefits, and we do this more efficiently and with an adaptive patch size.
-Additionally, our model is differentiable, meaning it can be used as a component in larger systems where accurate boundaries are necessary.
-Our work is complementary to recent approaches that leverage large-scale training to perform edge detection or segmentation by internalizing the patterns in a training set.
-Recent examples include Segment Anything [19] and combinations of strong self-supervised features with clustering and CRFs [5, 3, 20].
-These large-scale approaches rely on large training sets and high-capacity architectures that can internalize high-level cues like object and scene familiarity as well as low-level cues like spatial coherence.
-Our approach is different because it relies exclusively on low-level cues, has much less capacity, and uses much less training data.
-Despite this difference, we find that our model’s inherent preference for contour-like boundaries allows it to achieve comparable (or better) results, in particular when noise is high.
-This suggests that future work could benefit from incorporating some of our architectural elements into larger and higher-capacity systems, in particular to improve interpretability, efficiency, generalization or task-transfer.
+Early approaches to edge detection rely primarily on low-level or local information.
+Canny [3] and others (e.g.[5, 16, 10, 9]) use carefully chosen filters, and later ones like pB [15] and gPb [14] combine such filters using a handful of trainable parameters.
+These methods provide local estimates of straight-edge position and orientation and they can subsequently be filtered and joined using geometry-based processes such as edge-linking [3].
+However, they often struggle near corners and junctions due to the difficulty of designing filters for these more complicated structures.
+Structured edges [7] and other early learning-based methods [6, 13] are able to detect more-complicated structures, but they represent their local outputs using discrete, rasterized boundary maps that are not directly compatible with continuous boundary parameterizations or geometry-based linking.
+Recently, the field of junctions (FoJ) [25] introduced a way to detect more-complicated local structures while also representing them parametrically, without rasterization.
+By using its unrasterized representations in a customized non-convex optimization routine, the method can detect curves, corners and junctions with high precision and with unprecedented resilience to noise.
+Our model is strongly inspired by this work, and it improves upon it by introducing an enhanced family of parameterized geometric primitives (Figure 2) and by replacing its hand-designed objective and associated non-convex optimization routine with a learned model that is fast, feed-forward and differentiable.
+We also tackle the challenge of representing images that vary in scale across different areas.
+Unlike many low-level methods that assess image edges at a single, frequently manually chosen scale, our approach considers variations in scale throughout the image.
+The field of junctions, for example, requires setting a single patch size that junctions are optimized over, where larger patch sizes allow for better resilience to noise at the expense of memory and detail preservation.
+In contrast we equip each local region with a function to modulate its spatial extent.
+Our work is very different from the recent trend in computer vision of using deep networks to detect boundaries that are semantically meaningful. Examples include HED [27], EDTER [19], and UAED [28].
+These models are trained using annotated segmentation datasets like BSDS500 [2] and NYUDv2 [22], and they aim to identify boundaries between semantic objects (e.g., people, animals) while ignoring boundaries that are inside of these objects (e.g., stripes on a shirt).
+Their large capacities and their end-to-end training allow them to exploit not only low-level cues but also high-level cues such as object and scene familiarity.
+In contrast to this recent trend, we follow earlier work by focusing entirely on identifying boundaries from the bottom up, using low-level cues alone.
+Since our model is trained using simple synthetic data, it does not exploit object familiarity or learn to ignore intra-object edges.
+We find that this approach maintains some advantages: instead of producing a pixelized boundary map, our model produces a rich field of unrasterized primitives that provide better precision as well as information about local grouping.
+Moreover, despite not emphasizing top-down information, we find it performs better than deep models in the presence of high noise levels.
+This suggests that future work may benefit from incorporating some of our architectural elements into larger and higher-capacity systems, in particular to improve interpretability, efficiency, generalization and task-transfer.
 3 Representation
 Our system is depicted in Figure 1.
-It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with pixel-resolution tokens.
-Critically, each pixel’s token is explicitly required to encode a tuple of values representing a geometric primitive that specifies the geometric structure of the local boundaries surrounding the pixel.
-Because the tokens are forced through this geometric bottleneck, we refer to our model’s core mechanism as Boundary Attention.
-We instantiate our bottleneck using a learned linear mapping from the token dimension to a predefined lower-dimensional space of unrasterized boundary patterns that we call junction space.
-Our choice for this space is inspired by the Field of Junctions [33], but we use a modified parameterization (described below) that is differentiable across its entire domain.
+It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with stride-1 patch embeddings.
+Critically, each embedding explicitly encodes a tuple of values that specifies the geometric structure and spatial extent of the local boundaries within a patch.
+Because the embeddings are forced through this geometric bottleneck, we refer to our model’s core mechanism as boundary attention.
+We instantiate our bottleneck using a learned linear mapping from the embedding dimension to a predefined lower-dimensional space of unrasterized boundary patterns that we call junction space.
 Junction space has the benefit of specifying local boundary patterns without rasterization and thus with unlimited spatial precision.
 As depicted in Figure 2 and described in [25], it also has the benefit of including a large family of local boundary patterns, including uniformity (i.e., absence of boundaries), edges, bars, corners, T-junctions and Y-junctions.
+We also modulate the spatial extent of each junction by learning an associated windowing function, described in Section 3.1.
+To enable communication between patch embeddings, each of which corresponds to a local region of an image, we leverage the idea that each region of the image is described by multiple patches, and that these patches should agree on compatible representations of the global image structure.
+The ideal embedding for each patch should aptly represent its local geometry while agreeing with the structural claims of neighboring patches.
 The next section introduces our parameterization of junction space and some associated operators.
-Then Section 4 describes the architecture that we use to analyze an image into its field of junction values.
-Figure 2: Samples on a smooth trajectory through our space of geometric primitives (“junction space”), with one sample enlarged and accompanied by a visualization of its distance map on the right.
-Each sample comprises a vertex position $(u,v)$, orientation $\theta$, and angles $(\omega_{1},\omega_{2},\omega_{3})$ that partition the circle.
-The space includes all types of locally-uniform regions, edges, bars, corners, Y-junctions and T-junctions.
+Then Section 4 describes the architecture that we use to transform an image into its junction space representation.
 3.1 Boundary Primitives
@@
 Denote the $K$-channel input image by $\{\mathbf{f}[n]\}$, where $\mathbf{f}[n]\in\mathbb{R}^{K}$ is the vector image value at the discrete pixel grid index $n$.
-Our approach is to analyze the image into a field of dense, stride-1 overlapping local patches, each having a square support $\Omega_{n}(x)$ centered at the $n$th pixel.
-There are many ways to partition a local region $\Omega_{n}(x)$, and one can define parametric families of such partitions.
+Our approach is to analyze the image into a field of dense, stride-1 overlapping local patches, each having a square support $\Omega_{k}(x)$ for the $k$th patch centered at the $k$th pixel.
+There are many ways to partition a local region $\Omega_{k}(x)$, and one can define parametric families of partitions.
 For example the set of oriented lines provides a two-parameter family of partitions, with each member of the family separating the region into points that lie on one side of a line or the other.
@@
 We use the notation $\mathbf{g}=(\boldsymbol{u},\theta,\boldsymbol{\omega})$, where $\boldsymbol{u}=(u,v)\in\mathbb{R}^{2}$ is the vertex, $\theta\in\mathbb{S}^{1}$ is the orientation, and $\boldsymbol{\omega}=(\omega_{1},\omega_{2},\omega_{3})$ are barycentric coordinates (defined up to scale) for the three relative angles, ordered clockwise starting from $\theta$.
-Our convention is to express the vertex coordinates relative to the center of region $\Omega_{n}(x)$, and we denote that the vertex is free to move outside of this region.
+Our convention is to express the vertex coordinates relative to the center of region $\Omega_{k}(x)$, which is located at $c[k]$, and we denote that the vertex is free to move outside of this region.
 We also note that up to two angles $\omega_{j}$ can be zero.
-This all makes it possible to represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions).
-Fixing a value for $\mathbf{g}$ induces three binary-valued spatial support functions:
-$s_{nj}(x;\mathbf{g})\in\{0,1\},\ j=1,2,3,$ (1)
-that evaluate to $1$ for points $x$ that are in $\Omega_{n}(x)$ and in the $j$th wedge defined by $\mathbf{g}$; and 00 otherwise. It also induces an unsigned distance function:
-$d_{n}(x;\mathbf{g})\geq 0,$ (2)
-that represents the Euclidean distance from point $x$ to the nearest point in the boundary set defined by $\mathbf{g}$.
-Defining these as continuous functions allows specifying the local boundary structure with unlimited spatial resolution.
-The right of Figure 2 uses three colors to visualise the wedge supports of a junction $\mathbf{g}$, and it shows a quantized visualization of the associated distance function.
-Instead of fixing the size of every patch in the image to a hand-selected value that is the same throughout, we would like the size of every region $\Omega_{n}$ to adapt to the local geometry.
-We do this by equipping each patch with a parameterized local windowing function $w_{n}(x;\mathbf{p})\in[0,1]$.
-Specifically, we parameterize them as convex combinations of $W$ square window function coefficients $\mathbf{p}\in{\cal P}=\triangle^{W-1}$.
+This all makes it possible to represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions), and do so with unlimited spatial resolution.
+Fixing a value for $\mathbf{g}$ induces three wedge support functions, denoted by
+$s_{kj}(x;\mathbf{g})\in\{0,1\},\ j=1,2,3.$ (1)
+These evaluate to $1$ for points $x$ that are in $\Omega_{k}(x)$ and in the $j$th wedge defined by $\mathbf{g}$; and 00 otherwise. It also induces an unsigned distance function, denoted by
+$d_{k}(x;\mathbf{g})\geq 0,$ (2)
+which represents the Euclidean distance from point $x$ to the nearest point in the boundary set defined by $\mathbf{g}$.
+Figure 2 uses three colors to visualise the wedge supports $s_{kj}$ of a junction $\mathbf{g}$, and Figure 3 shows the unsigned distance functions for $3\times 3$ grids of junction parameters.
+Analytic expressions for them are included in the supplement.
+In order to enable the size of each region $\Omega_{n}$ to be able to adapt to the local geometry and noise conditions, we equip each one with a parameterized local windowing function $w_{k}(x;\mathbf{p})\in[0,1]$, with parameters $\mathbf{p}\in{\cal P}=\triangle^{W-1}$ that are the coefficients of a convex combination of $W$ square pillbox functions.
 That is,
-$w_{n}(x;\mathbf{p})=\sum_{i=1}^{W}p_{i}\mathbf{1}[\|x-c[n]\|_{\infty}\leq D_{i}],$ (3)
-where $\|\cdot\|_{\infty}$ is the $\ell^{\infty}$-norm, and $\mathbf{1}[\cdot]$ is the indicator function that returns $1$ if the argument true; and 00 otherwise.
-In our experiments we use $W=3$ and diameters $\mathbf{D}=(3,9,17)$.
-Figure 3: Visualizing our model’s output.
-Top row, left to right: (i) input image; (ii) boundary-aware smoothing; and (iii) boundary map with attention maps at four query points.
-The two bottom rows show regions $\Omega_{n}$ unfolded from within the two white windows indicated above.
-Bottom rows, left to right: (i) window functions; (ii) boundaries; and (iii) support functions colored by their gathered wedge features.
-See text for details.
+$w_{k}(x;\mathbf{p})=\sum_{i=1}^{W}p_{i}\mathbf{1}[\|x-c[k]\|_{\infty}\leq D_{i}],$ (3)
+[…]
+Figure 3: Example of our model’s output, with examples from two different regions.
+Top row: Some of each region’s overlapping input patches, their corresponding outputs (visualized in the style of Figure 2), and three types of per-patch attributes that the outputs imply: unsigned distance; boundaries; and gathered wedge colors.
+Bottom row: Four types of global maps that are implied by accumulating values from the output field and rendered patches.
+See the supplement for interactive visualizations.
+3.2 Gather and Slice Operators
+Our network operates by refining the field $\{(\mathbf{g}^{t}[k],\mathbf{p}^{t}[k])\}$ over a fixed sequence of steps $t=1,...,T$.
+It uses two operators that we define here and depict in the right of Figure 4 to facilitate intra-patch communication between pixels within a single patch, and inter-patch communication across different patches representing overlapping image regions.
+The first operator is a patch-wise gather operator, in which each wedge of each patch computes the weighted average of the image values it contains (recall that $c[n]$ are the $n$th pixel’s coordinates, and wedges are indexed by $j$):
+$\mathbf{f}_{kj}=\frac{\sum_{n}\mathbf{f}[n]w_{k}(c[n];\mathbf{p}[k])s_{kj}(c[n];\mathbf{g}[k])}{\sum_{n}w_{k}(c[n];\mathbf{p}[k])s_{kj}(c[n];\mathbf{g}[k])}.$ (4)
+The second operation is a global pixel-wise slice operation, where each pixel computes the means and variances, over all regions that contain it, of the per-region distance maps $d_{k}(x;\mathbf{g}[k])$ and gathered wedge features $\mathbf{f}_{kj}$.
+The expressions for the means are:
+$\bar{d}[n]$ $=\frac{\sum_{k}w_{k}(c[n];\mathbf{p}[k])d_{k}(c[n];\mathbf{g}[k])}{\sum_{k}w_{k}(c[n];\mathbf{p}[k])},$ (5)
+$\bar{\mathbf{f}}[n]$ $=\frac{\sum_{k}w_{k}(c[n];\mathbf{p}[k])\sum_{j}\mathbf{f}_{kj}s_{kj}(c[n];\mathbf{g}[k])}{\sum_{k}w_{k}(c[n];\mathbf{p}[k])\sum_{j}s_{kj}(c[n];\mathbf{g}[k])}.$ (6)
+Note that only patches that contain $c[n]$ contribute to the sum, i.e., the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$.
+The expressions for the pixel-wise distance map variance $\nu_{d}[n]$, and for the feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are defined accordingly and included in the supplement.
+Intuitively, slicing $\mathbf{f}_{kj}$ can be thought of globally accumulating information from each patch’s junction representation.
+3.3 Visualizing Output
+At our network’s output, we are provided with a field of tuples representing the junction and window parameters $\{(\mathbf{g}[k],\mathbf{p}[k]\}$ for all stride-1 patches of the input image.
+Though these outputs are defined continuously, we can visualize them in Figure 3 by rendering the windowing functions $w_{k}(x;\mathbf{p})$ (second column), and the binary junction wedge supports $s_{kj}(x;\mathbf{g})$ colored purple, orange, and blue (third column), along with the input image patches (first column).
+To the left, we show the input image and the regions from which the patches were extracted from.
+Additionally, we can use the learned junction parameters to render unsigned distance patches $d_{k}(x;\mathbf{g})$ (fourth column), boundary patches $b_{k}(x;\mathbf{g})$ (fifth column), and the binary junction wedges $s_{kj}(x;\mathbf{g})$ colored with their respective gathered colors (sixth column).
+All of the patches are defined continuously, so we can render them at any resolution.
+We expect the shapes of junction boundaries in overlapping regions to agree, so that the variances $\nu_{d}[n],\nu_{f}[n]$ are small at every pixel.
+Then, the fields of means $\{\bar{d}[n]\},\{\bar{\mathbf{f}}[n]\}$ can be interpreted, respectively, as a global unsigned distance map for the image boundaries (bottom row, fourth column) and a boundary-aware smoothing (bottom row, last column) of its input channel values.
+Figure 3 shows an example (bottom row, fifth column), where we visualize the zero-set of the global unsigned distance map—we call this the global boundaries—by gathering boundary patches $b_{k}$ defined as:
+$b_{k}(x;\mathbf{g})=\left(1+(d_{k}(x;\mathbf{g})/\eta)^{2}\right)^{-1},$ (7)
+setting $\eta=0.3$.
+The output contains much more information than these fields of means.
+For any query pixel $n$, we can probe the wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ that contain it, thereby obtaining a spatial attention map $a_{k}(x)$ that surrounds the query pixel.
+This is the boundary-aware spatial kernel that turns a neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the value $\bar{\mathbf{f}}[n]$:
+$\bar{\mathbf{f}}[n]=\sum_{k}a_{n}(c[k])\mathbf{f}[k].$ (8)
+The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6, and its maximum diameter is twice that of $\Omega(x)$.
+The spatial affinity maps for two probed points are shown in the leftmost image of the bottom row of Figure 3.
+Like slicing, probing the affinity map around a pixel is a form of global accumulation that takes values defined per-patch and synthesizes them into a global representation.
 Figure 4: Model Architecture.
@@
 There are $207$k trainable parameters in total.
-3.2 Gather and Slice Operators
-Our network operates by refining the field $\{(\mathbf{g}^{t}[n],\mathbf{p}^{t}[n])\}$ over a fixed sequence of steps $t=1,...,T$.
-It uses two operators that we define here and depict in the right of Figure 4.
-The first operator is a patch-wise gather operator, in which each wedge of each region computes the weighted average of the image values it contains (recall that $c[n]$ are $n$th pixel’s coordinates):
-$\mathbf{f}_{kj}=\frac{\sum_{n}\mathbf{f}[n]w_{k}(c[n];\mathbf{p}[n])s_{kj}(c[n];\mathbf{g}[n])}{\sum_{n}w_{k}(c[n];\mathbf{p}[n])s_{kj}(c[n];\mathbf{g}[n])}.$ (4)
-The second operation is a pixel-wise slice operation, where each pixel computes the means and variances, over all regions that contain it, of the per-region distance maps $d_{n}(x;\mathbf{g}[n])$ and gathered wedge features $\mathbf{f}_{kj}$.
-The expressions for the means are:
-$\bar{d}[n]$ $=\frac{\sum_{k}w_{k}(c[n];\mathbf{p}[k])d_{k}(c[n];\mathbf{g}[k])}{\sum_{k}w_{k}(c[n];\mathbf{p}[k])},$ (5)
-$\bar{\mathbf{f}}[n]$ $=\frac{\sum_{k}w_{k}(c[n];\mathbf{p}[k])\sum_{j}\mathbf{f}_{kj}s_{kj}(c[n];\mathbf{g}[k])}{\sum_{k}w_{k}(c[n];\mathbf{p}[k])\sum_{j}s_{kj}(c[n];\mathbf{g}[k])}.$ (6)
-Note that the only contributions to the sums over patches $k$ are from patches that contain $c[n]$, i.e., the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$.
-The expressions for the pixel-wise distance map variance $\nu_{d}[n]$, and for the feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are defined accordingly and included in the supplement.
-3.3 Visualizing Output
-At our network’s output, we expect the shapes of junction boundaries in overlapping regions to agree, so that the variances $\nu_{d}[n],\nu_{f}[n]$ are small at every pixel.
-Then, the fields of means $\{\bar{d}[n]\},\{\bar{\mathbf{f}}[n]\}$ can be interpreted, respectively, as a global unsigned distance map for the image boundaries and a boundary-aware smoothing of its input channel values.
-Figure 3 shows an example, where we visualize the zero-set of the global unsigned distance map—we call this the global boundary map—by applying the nonlinearity:
-$\bar{b}_{\eta}[n]=\left(1+(\bar{d}[n]/\eta)^{2}\right)^{-1},$ (7)
-setting $\eta=0.7$.
-The output contains much more information than these fields of means.
-For any query pixel $n$, we can probe the wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ that contain it, thereby obtaining a spatial attention map $a_{n}(x)$ that surrounds the query pixel.
-This is the boundary-aware spatial kernel that turns a neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the value $\bar{\mathbf{f}}[n]$:
-$\bar{\mathbf{f}}[n]=\sum_{k}a_{n}(c[k])\mathbf{f}[k].$ (8)
-The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6, and its maximum diameter is twice that of $\Omega(x)$. Some examples are shown in the top-right of Figure 3.
-Additionally, as shown in the bottom two rows of Figure 3, we can unfold any portion of the output field into the overlapping regions it contains.
-The figure shows unfolded regions $\{\Omega_{k}(x)\}$ within two windows of the output field.
-For each unfolded set we visualize: the windowing functions $w_{k}(x;\mathbf{p})$; the regional boundaries $b_{k}(x;\mathbf{g})$ obtained by applying the nonlinearity in Equation 7 to the local distance functions $d_{k}(x;\mathbf{g})$; and the supports $s_{kj}(x;\mathbf{g})$ colored according to the wedge features $\mathbf{f}_{kj}$ that they gather from the input image.
-In the top example, the neighborhood is homogeneous, so the windowing functions (via $\mathbf{p}[n]$) have large supports, and there are few boundaries (via $\mathbf{g}[n]$).
-In the bottom example there is fine-scale geometry.
-The windowing functions are narrow and the primitives agree on a Y-junction.
 4 Network Architecture
-We design our network to iteratively refine the fields $\{(\mathbf{g}^{t}[n],\mathbf{p}^{t}[n])\}$.
-We do so by embedding each field element using a higher dimensional representation, $\boldsymbol{\gamma}^{t}[n]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[n]\in\mathbb{R}^{D_{\pi}}$ respectively, which can be updated via dot-product attention.
-In practice, we use $D_{\gamma}=64$ and $D_{\pi}=8$, which provides the network with enough capacity to learn meaningful hidden states.
-We learn simple linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$ which are used for the gather and slice operations.
-Importantly, all of our network’s elements are invariant to discrete spatial shifts of the image, operating either on individual locations $n$ or on small neighborhoods of locations.
-See Figure 4 and additional details in the supplement.
-Given an input image, the network first applies a neighborhood MLP-mixer, which is a modified variation of MLP-Mixer [32] where we replace the linear operations with convolutions of kernel size $3$.
+We seek a differential architecture that can effectively initialize and then refine the fields $\{(\mathbf{g}^{t}[k],\mathbf{p}^{t}[k])\}$ using a manageable number of iterations.
+This bears similarities to the edge localization and edge-linking steps of early edge detectors like Canny’s, but with a more sophisticated local boundary model, and with mechanisms for spatial consistency that are learned instead of designed.
+It is also analogous to the original field of junctions algorithm [25], which uses coordinate descent for initialization and iterations of gradient descent for refinement.
+We want to replace both steps with something that is differentiable, faster, and able to scale to larger images.
+After considering a variety of alternatives, we settled on a particular family of architectures based on dot-product self-attention; and within this family, we performed an extensive set of ablations to determine the best performing model.
+We describe our final model here, and we provide the ablation details in Section 6.
+Importantly, all of our model’s components are invariant to discrete spatial shifts of the image, operating either on individual locations $k$, or on small neighborhoods of locations with spatially-shared parameters.
+This means that our model can be trained on small images and then deployed on much larger ones.
+Also, our model has only $207$k learnable parameters, making it orders of magnitude smaller than most deep semantic boundary detectors.
+As a point of reference, the full EDTER model has 109M parameters [19].
+Our model is depicted in Figure 4.
+It implicitly uses higher-dimensional embeddings of the field elements, $\boldsymbol{\gamma}^{t}[k]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[k]\in\mathbb{R}^{D_{\pi}}$, which can be decoded at any iteration using the learned linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$.
+Our final model uses $D_{\gamma}=64$ and $D_{\pi}=8$ which, as shown in the supplement, provides enough capacity to learn a smooth latent representation of junction space.
+Use separate embeddings for the junction and windowing fields provides a disentangled representation of both.
+Given an input image, the network first applies a “neighborhood MLP-mixer”, which is a modified MLP-Mixer [24] that replaces the global spatial operations with convolutions of kernel size $3$.
 The other change is that we map the input pixels to the hidden state size with a pixel-wise linear mapping rather than taking patches of the input.
-This block, which we denote the “neighborhood MLP-mixer” transforms the input into an initial hidden state, which is then processed through eight iterations of our boundary attention to further refine the hidden state.
-The eight iterations of refinement are broken into two Boundary Attention Blocks with independent weights.
+This block transforms the input image into a pixel-resolution map with ${D_{\gamma}}$ channels.
+We denote this by $\gamma_{0}[n]$ and refer to it as the initial “hidden state”.
+This hidden state is then refined by a sequence of eight boundary attention iterations, which we describe next.
+(See the supplement for a visualization of the decoded hidden states as they evolve.)
+The eight iterations of refinement are broken into two boundary attention blocks, each with learned weights.
 To process our input, we first add in a linear projection of the initial hidden state.
-This is essentially a skip connection that allows our network to retain information from the input pixels at later stages of processing.
+This is essentially a skip connection that allows our network to retain information from the initial hidden state estimate at later stages of processing.
 Next, we copy our hidden state into two identical pieces.
-We concatenate a dimension 8 learned windowing token to one of the copies and the current estimate of the smoothed global features to the other.
+We concatenate a dimension 8 learned windowing embedding to one of the copies and the input image plus the current estimate of the smoothed global features to the other.
 We then do neighborhood cross-attention: each pixel in the first copy does two iterations of cross attention with a size 11 patch of the second copy.
-We add a learned $11\times 1$ 11 positional encoding to the patch, which allows our network to access relative positioning, even if global position cues are absent.
+We add a learned $11\times 1$ 11 positional encoding to the patch, which allows our network to access relative positioning, even though global position cues are absent.
 We follow each self attention layer with a small MLP.
 To transform our output or intermediary hidden state into junction space and render output images, we use a simple linear mapping.
-We separate the windowing token (the last 8 dimensions) from the hidden state (the first 64 dimensions) and project each through a linear layer.
-We map the hidden states to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$.
+We separate the windowing embedding (the last 8 dimensions) from the junction embedding (the first 64 dimensions) and project each through a linear layer.
+We map the hidden state to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$.
 These serve as the inputs to our gather and slice operators.
-Overall, our network has $2.07\cdot 10^{5}$ learnable parameters, making it orders of magnitude smaller than most learned boundary detectors.
-For contrast, EDTER [28] has 109 million parameters for its full model.
+To extract $\mathbf{p}$ from the windowing embedding, we linearly project the windowing embedding to a length 3 vector, which we use as coefficients in a weighted sum of three square pillbox functions with widths $3$, $9$, and $17$.
+This implementation of the windowing function ensures spatial overlap between neighboring patches by limiting the minimum patch size to $3$.
 4.1 Training
 We train our network in two stages: we begin by training the neighborhood MLP-mixer with the first Boundary Attention Block, and then we add the second block and retrain end-to-end.
+Because junction space is non-convex and difficult to optimize over, this two step training procedure helps the network converge.
 We apply our loss to the two final iterations of our network, so for the first stage of training we apply a loss to iterations 3 and 4, and for end-to-end refinement we apply a loss to iterations 7 and 8.
-We weigh the final loss three times higher than the second to last loss, which encourages the network to allocate capacity to producing high quality outputs, while ensuring that gradient information is shared across the network.
+We weigh the final loss three times higher than the second to last loss, which encourages the network to allocate capacity to producing high quality outputs, while still providing supervision to intermediate junction estimates.
 We train our method using a combination of four global losses applied to global (i.e.
-averaged) fields, and two patch-wise losses applied to individual patches.
+sliced) fields, and two patch-wise losses applied to individual patches.
 The first two losses are supervision losses penalizing mismatches between our network’s predictions and the ground truth feature and boundary maps:
@@
 where $\mathbf{f}_{\text{GT}}$ and $d_{\text{GT}}$ are the ground truth features and distance maps, respectively, and $\alpha[n]$ is a pixel importance function defined as:
-$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+C,$ (11)
+$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+(1-C),$ (11)
 with $\beta$ and $C$ controlling how much weight to give pixels near boundaries.
-We set $\beta=0.1$, $\delta=1$, and increase $C$ throughout training to give more weight to non-boundary locations.
+We set $\beta=0.1$, $\delta=1$, and anneal $C$ to give more weight to non-boundary locations later in training.
 We also tested a more involved pixel importance mask that gave extra weight to visible vertices and intersections, the details for which can be found in the supplement.
@@
 and we set $\delta^{\prime}=1$. These per-patch losses gives the network a more direct signal on how to adjust its weights than purely global losses, which average over multiple patches.
-Figure 5: The ODS F-score of our method at different noise levels, compared with the Field of Junctions (FOJ) [33] run using patch sizes $9\times 9$, $17\times 17$, and $25\times 25$, and EDTER [28], with and without preprocessing by BM3D [9].
-The bottom insets show example patches at representative PSNR values.
+Figure 5: Left: ODS F-score for our method and multiple baselines at different noise levels computed on noisy synthetic data.
+The bottom inset show example patches at representative PSNR values.
+Our method outperforms all baselines at low noise and is better or competitive with other techniques at high noise.
+Right: Comparing the F-score for different techniques with their runtime.
+Our method has the best average F-score while also being much faster than the second best method Field of Junctions.
+Figure 6: Left: Qualitative comparison of our model versus other methods for noisy crops from the ELD dataset [26].
+Our method provides more detail than other techniques at low noise and is more robust to high levels of noise.
+Right: Repeatability of the estimated boundaries of crops from the ELD dataset over increasing noise.
+Over multiple levels of real sensor noise, our method is the most consistent in predicting edges.
 5 Experiments
@@
 To those crops we add Gaussian and/or Perlin noise [17], and with probability $0.1$ we average over the color channels to produce grayscale inputs.
-Our dataset contains $10^{5}$ images, $90\%$ of which are used for training, and the rest for testing.
-For training and optimization details, refer to the supplement.
-Performance w.r.t. noise levels.
-Figure 5 shows the comparison of our method and baseline approaches, the Field of Junctions [33] and EDTER [28] under different noise levels.
-We test the Field of Junctions with different patch sizes ($9\times 9$, $17\times 17$, and $25\times 25$) and EDTER with/without preprocessing using optimally-tuned denoising using BM3D [9].
+Our dataset contains $10^{5}$ images, $90\%$ of which are used for training, and the rest for validation.
+For a visualization of training data and training and optimization details, refer to the supplement.
+Baselines.
+Since we focus on low-level edge detection, we primarily compare against other techniques with a similar focus and that also do not train on large semantic datasets: Canny [3], gPb [14], Structured Edges (SE) [7], and Field of Junctions (FoJ) [25].
+However, since most recent work tends to focus on semantic edge detection, in the interest of being comprehensive, we also provide a comparison to several modern works: UAED [28], EDTER [19], and HED [27].
+We note that this is not a clean comparison due to the difference in tasks – predicting edges with low-level cues vs predicting edges with semantic meaning.
+Quantitative results and inference times across different noise levels.
+For quantiative evaluation, our focus on low-level cues means that we cannot use semantic edge detection benchmarks like BSDS50 [2] since our method and other low-level baselines will both predict edges that are not semantically meaningful and fail to predict edges that are semantically meaningful but lack contrast.
+We instead rely on synthetic data where the ground truth edges can be determined with perfect precision and any level of noise can be added.
+In this experiment, we use geometric objects and per-pixel additive Gaussian noise with a random variance.
+Figure S3 compares the performance and inference time of our method and baselines under different noise levels.
 The tuneable parameters for Field of Junctions were chosen to maximize its performance on noisy images with $17\times 17$ patches.
-Our method outperforms all baselines at lower noise levels and is competitive with the Field of Junctions at higher noise levels, while being orders of magnitude faster than it (see Table 1).
-Figure 6: Our network can reliably locate boundaries with sub-pixel precision, due in part to our parametric representation for boundaries.
-We calculate the ODS F-Score of synthetic images as we increase the pixel matching distance threshold (how close a pixel must be to a ground truth pixel to be counted) from 0 to 1.5 and vary the PSNR of the images from $\infty$ to 8.0.
-Each line represents the performance at a noise level, with lighter-colored lines corresponding to noisier inputs.
-X-axis values below 1.0, denoted by the vertical dotted line, represent sub-pixel precision.
-Figure 7: One of the advantages of our parametric form for boundaries is that we can trivially upsample our output boundary maps by striding the predicted junction parameters, and scaling the patch size accordingly.
-This results in clean boundaries even when our output boundary images are upsampled from $125\times 125$ (left) to $500\times 500$ (right).
-Figure 8: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$.
-Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$.
-The embedding has learned to be smooth and have an intuitive zero.
-Sub-pixel precision.
-To measure sub-pixel precision, we render $500\times 500$ images containing pairs of overlapping circles and triangles.
-This gives us a precise binary boundary map for that resolution.
-We downsample the images to $125\times 125$, and add varying amounts of Gaussian noise.
-These noisy, downsampled images serve as the inputs.
-To evaluate the predicted boundaries, we upsample the outputs back to $500\times 500$ pixels.
-For EDTER, we use bilinear interpolation along with edge thinning to produce the upsampled boundaries.
-In our case, a byproduct of our parametric form for boundaries is that an intuitive method for upsampling boundaries naturally arises: by increasing the patch stride to $4$, and rendering the patches as $4$ times as large—68×$\times 68$ 68 pixels—we can create an upsampled version of our image while retaining our method’s accuracy in boundary localization.
-We visualize this process in Figure 7.
-We evaluate the upsampled outputs on the original $500\times 500$ binary ground truth maps and vary the minimum distance threshold of the evaluation metric, which denotes how close a prediction must be to the ground truth to be used.
-The results are shown in Figure 6.
-Our F-score remains high across all matching thresholds, even when the inputs contain additive Gaussian noise.
-Figure 9: Evolution of boundaries during iterations, in reading order.
-Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
-Figure 10: Our model generalizes well to real images, and can find accurate boundaries in natural images.
-Top left: The input image, a low light image from the SIDD [1] dataset.
-Top right: Our network’s predicted boundaries.
-Bottom left: The output mean feature map.
-Bottom right: The predicted distance map.
-Linear interpolation in junction space.
-Surprisingly, we note that our network learns a spatially smooth manifold of junctions in its hidden state.
-Figure 8 visualizes some properties of its learned embedding of junction space.
-We generate equally-spaced samples $\boldsymbol{\gamma}_{i}\in\mathbb{R}^{D_{\boldsymbol{\gamma}}}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to 00 and then to a particular $\boldsymbol{\gamma}_{b}$.
-We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center.
-For visual comparison, we show an analogous interpolation from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$ by using a linear geometric interpolation in junction space ${\cal G}$, the expressions for which are in the supplement.
-Evolution of the outputs over time.
-Figure 9 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement.
-Specifically, we visualize the result of slicing similar to Equation 5 but with the regional distance functions $d_{k}$ replaced by their nonlinear counterparts $b_{k}$.
-We see that early iterations are exploratory and unstructured, and that later iterations reach agreement.
-Results on real images.
-Figures LABEL:fig:real_results and 10 show the qualitative results on real images.
-Despite being trained on synthetic data, our method can outperform existing state-of-the-art methods at multiple levels of real sensor noise present in ELD [34].
+Notably our method’s adaptive windowing function gives it at edge compared to the Field of Junctions at low noise, enabling it to capture finer details, with only slightly worse performance under extreme noise conditions.
+Our method is also orders of magnitude faster than FoJ as shown in the plot to the right.
+Qualitative results on real images.
+As shown in Figure 6, despite being trained on synthetic data, our method can outperform existing state-of-the-art methods at multiple levels of real sensor noise present in ELD [26].
 Our method produces crisp and well-defined boundaries despite high levels of noise.
-Table 1: Inference time (in seconds) of EDTER [28] (without preprocessing by BM3D [9]), and Field of Junctions (FoJ) [33] run at various patch sizes, compared with our method.
-We compare both with the original (FoJ) implementation as well as our JAX reimplementation (JAX-FoJ).
-Unlike FoJ [33], our method does not require striding for fitting moderately-sized images in memory (runtimes reported with $*{}^{*}$ did require striding).
-The average run times were calculated on an Nvidia A100 GPU.
-$125\times 125$ $320\times 320$
-EDTER [28] $0.130$ $0.130$
-FoJ [33]  $9\!\!\times\!\!9$ $36.8$ $206$
- $17\!\!\times\!\!17$ $88.0$ $76.0^{*}$ (stride-$3$)
- $25\!\!\times\!\!25$ $162$ $57.6^{*}$ (stride-$5$)
+Figure 7 shows that our method makes reasonable boundary estimates for even complex in-the-wild images.
+Our supplement contains additional examples.
+Figure 7: Our model can find compelling boundaries in natural in-the-wild images, despite training on simple synthetic data.
+Left image from the SIDD [1] dataset; right image taken with a phone camera.
+Repeatability on real images.
+We quantify each model’s noise resilience on real low-light images by measuring the repeatability of its boundaries over a collection of scenes from the ELD dataset.
+For each scene, we run the model on the lowest-noise image and then measure, via ODS F-score, how much its predictions change with increasing noise level.
+Figure 6 shows the averages of these scores across the collection.
+Our model provides more consistent results for increasingly noisy images than other methods.
+6 Ablations
+[…]
+table:ablation1 shows what happens when the initial per-patch Neighborhood MLP-mixer [NM] is used alone, versus combining it with one boundary attention block [BA-1] or two such blocks [BA-2].
+Tying the attention weights across iterations provides a slight advantage at higher noise levels, with only a slight penalty in accuracy at lower noise.
+Our final model is boxed.
+It provides the best overall performance, accepting slightly lower accuracy at low noise in exchange for better accuracy at higher noise levels.
 […]
@@
 […]
-Ours $0.0823$ $0.678$
-Inference time.
-Table 1 compares the running time for different methods at two resolutions.
-We denote several runs for Field of Junctions where we had to increase the patch stride (optimize junctions at strided intervals) to avoid running out of memory.
-Notably, we are orders of magnitude faster than Field of Junctions, despite achieving similar performance.
-Additionally, because our network can handle inputs of any size, the run time depends on the input shape, in contrast with EDTER that has a maximal input size of $320\times 320$ inputs and constant run time.
+N/A $0.355$ $0.189$ $0.179$
+$0.837$ $0.582$ $0.293$
+0.874 $0.658$ $0.327$
+$0.872$ 0.673 0.348
+Table 1: Ablations: impact of component combinations and weight-tying on F-score (higher is better). Our reported model is boxed.
+[…]
+Iterations 2 $0.857$ $0.611$ $0.307$
+3 $0.867$ $0.647$ $0.329$
+4* 0.872 0.673 0.348
+5 $0.871$ $0.663$ $0.337$
+Neighborhood $3\times 3$ $0.835$ $0.531$ $0.240$
+$7\times 7$ $0.869$ $0.637$ $0.313$
+$9\times 9$ $0.871$ $0.658$ $0.325$
+$11\times 1$ 11* 0.872 0.673 0.348
+$13\times 13$ $0.871$ $0.655$ $0.324$
+MLP Input Constant $0.870$ $0.650$ $0.317$
+Input features $0.868$ $0.657$ $0.316$
+Avg. features* 0.872 0.673 0.348
+Loss Uniform $0.844$ $0.629$ $0.313$
+Weighted* 0.872 0.673 0.348
+Windowing Fixed - $3\times 3$ $0.834$ $0.546$ $0.205$
+Fixed - $9\times 9$ 0.872 $0.662$ $0.328$
+Fixed - $17\times 17$ $0.859$ $0.662$ $0.328$
+Inferred* 0.872 0.673 0.348
+Table 2: Ablations: variations of our reported model.
+We compare the number of iterations of boundary attention, the attention neighborhood size, the features concatenated to the hidden state as input to the MLP, the impact of a spatially weighted loss, and a fixed versus learned windowing function.
+Asterisks indicate our reported choices.
 7 Conclusion
-We have introduced a differentiable model that explicitly reasons about geometric primitives such as edges, corners, junctions, and regions of uniform appearance, by using boundary attention.
-Despite being trained on simple geometric synthetic data, our method generalizes to natural images, and predicts clean boundaries, even when the images it is applied to are significantly corrupted by noise.
-Furthermore, the parametric form of our model elicits a natural formulation for boundaries with subpixel precision, and its bottom-up patch-wise approach enables it to process any image at its native resolution and aspect ratio.
-Supplemental Material
-Boundary Attention: Learning to Find Faint Boundaries at Any Resolution
-1 . The space of $M$-junctions
+We have introduced a differentiable model that uses boundary attention to explicitly reason about geometric primitives such as edges, corners, junctions, and regions of uniform appearance in images.
+Our bottom-up, feedforward network can encode images of any resolution and aspect ratio into a field of geometric primitives that describe the local structure within every patch.
+Despite being trained on simple geometric synthetic data, our method can consistently localize boundaries under high levels of noise better than state-of-the-art deep networks, while achieving similar performance to recent optimization-based approaches in a fraction of the time.
+Supplemental Material for Boundary Attention: Learning to Localize Boundaries under High Noise
+S1 The space of $M$-junctions
 Here we provide the expressions for the support functions $s_{j}(x;\mathbf{g})$ and the unsigned distance function $d(x;\mathbf{g})$ from Section 3 of the main paper.
@@
 This captures the fact that $M$-junction families are nested for increasing $M$.
-Figure 11: Anatomy of an $M$-junction $\mathbf{g}=(\mathbf{u},\theta,\boldsymbol{\omega})$ with $M=3$.
+Fig.
+S1: Anatomy of an $M$-junction $\mathbf{g}=(\mathbf{u},\theta,\boldsymbol{\omega})$ with $M=3$.
 Left: Boundary directions $\phi_{j}$ and central directions $\psi_{j}$ are determined directly from relative angles $\boldsymbol{\omega}$ and orientation $\theta$ (which is equal to $\phi_{1}$).
@@
 Right: Associated boundary function $b_{\eta}(x;\mathbf{g})$ using $\eta=0.7$.
-As shown in Figure 11, other geometric features of a junction can be directly derived from the orientation and angles. The central directions $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{M})$ are
+As shown in Figure S1, other geometric features of a junction can be directly derived from the orientation and angles. The central directions $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{M})$ are
 $\psi_{j}=\theta+\frac{\tilde{\omega}_{j}}{2}+\sum_{k=1}^{j-1}\tilde{\omega}_{k},\quad j\in\{1\ldots M\},$ (18)
@@
 $b_{\eta}(x;\mathbf{g})=\left(1+(d(x;\mathbf{g})/\eta)^{2}\right)^{-1}.$ (23)
-Figure 11 shows an example of a junction’s distance function and its associated boundary function with $\eta=0.7$.
-Interpolation
+Figure S1 shows an example of a junction’s distance function and its associated boundary function with $\eta=0.7$.
+S1.1  Interpolation
 Another advantage of the present parameterization compared to that of the original [25] is that it is a simply-connected topological space and so allows for defining mechanisms for smoothly interpolating between any two junctions $\mathbf{g}$ and $\mathbf{g}^{\prime}$.
@@
 The bottom row of Figure 9 in the main paper visualizes a set of samples from smooth trajectories in junction space using this mechanism.
-2 . Training Data
-Figure 12: Columns 1 to 5: Examples of the synthetic data used to train our model using supervision with ground-truth boundaries.
+S2 Training Data
+Fig.
+S2: Columns 1 to 5: Examples of the synthetic data used to train our model using supervision with ground-truth boundaries.
 Column 6: Rendered distance maps corresponding to column 5.
@@
 Each noiseless image has an unrasterized, vector-graphics representation of its shapes and colors, which specify the clean image and exact boundary-distance map with unlimited resolution.
-We find that we can train our model to a useful state using purely synthetic data, examples of which are depicted in Figure 12.
+We find that we can train our model to a useful state using purely synthetic data, examples of which are depicted in Figure S2.
 In fact, we find it sufficient to use very simple synthetic data that consists of only two basic shapes—circles and triangles—because these can already produce a diverse set of local edges, thin bars, curves, corners, and junctions, in addition to uniform regions.
@@
 Since zero-mean noise can at times result in values below 0 or above the maximum magnitude threshold, we truncate any pixels outside of that range.
-3 . Model Details
+S3 Model Details
 Our model is designed to be purely local and bottom up, with all of its compositional elements operating on spatial neighborhoods in a manner that is invariant to discrete spatial shifts of an image.
@@
 We implement our model in JAX, and we will publicly share our model code and its pretrained weights upon publication.
-S3.1 . Neighborhood MLP-Mixer
+S3.1  Neighborhood MLP-Mixer
 Our neighborhood MLP-mixer is a shift invariant, patch-based network inspired by MLP-mixer [24].
@@
 To handle border effects in our neighborhood MLP-mixer, we apply zero-padding after the initial projection from $\mathbb{R}^{3}$ to $\mathbb{R}^{64}$, and then we crop to the input image size after the second neighborhood mixing block to remove features that correspond to patches without full coverage, i.e., patches that contain pixels outside of the original image.
-S3.2 . Neighborhood Cross-attention
+S3.2  Neighborhood Cross-attention
 The neighborhood cross-attention block similarly enforces shift-invariance and weight sharing across spatial neighborhoods.
@@
 To handle border effects, we zero-pad the key and value tokens so that every query attends to an $11\times 1$ 11 neighborhood, and then zero-out any attention weights involving zero-padded tokens.
-S3.3 . Training details
+S3.3  Training Details
 We pretrain the neighborhood MLP-mixer and the first boundary attention block on a simplified variation of our Kaleidoshapes dataset, where each image is $100\times 100$ and contains a single triangle and circle with additive zero-mean Gaussian noise.
@@
 The final stage of training adds a second boundary attention block with weights that are initialized using a copy of the pretrained weights of the first boundary attention block.
-We use $100,000$ crops of size $125\times 125$ from our Kaleidoshape images (10% withheld for testing) and the full set of losses; and we optimize all of the model’s parameters, including those of the neighborhood MLP-mixer and the first boundary attention block.
+We use $100k$ crops of size $125\times 125$ from our Kaleidoshape images (10% withheld for testing) and the full set of losses; and we optimize all of the model’s parameters, including those of the neighborhood MLP-mixer and the first boundary attention block.
 Like in pretraining, we add a spatial importance that prioritizes region containing a corner (i.e., a visible triangle vertex) or a junction (i.e., a visible intersection between the boundaries of any two shapes).
 Our trained weights will be released upon publication.
-4 . Qualitative Behavior for Natural Images
-In Figures 14 and 13, we show how the model behaves on noiseless natural images that contain texture and recognizable objects.
-In particular, Figure 13 emphasizes how the boundary maps produced by our model qualitatively differ from those of many classical bottom-up edge-detectors and also from those of learned, end-to-end models that have been trained to match human annotations.
+S4 Model Behavior
+S4.1  Generalization to Unseen Shapes
+During training, our model observes only circles and triangles of varying sizes and colors.
+We measure how well our model generalizes to unseen shapes by creating a new synthetic dataset containing rectangles, polygons, and ellipses of various sizes, colors, and rotations.
+We evaluate our model against low-level methods as well as methods trained on semantic data, and note the results here closely align with the performance trends on kaleidoshapes (Fig.
+5), suggesting that our bottom-up model can generalize to unseen geometries.
+Fig. S3: We measure how well our model can generalize to shapes outside of the training distribution by evaluating our model and others on noisy synthetic data composed exclusively of unseen shapes.
+S4.2  Linear Interpolation in Junction Space
+Surprisingly, we note that our network learns a spatially smooth manifold of junctions in its hidden state.
+Figure S4 visualizes some properties of its learned embedding of junction space.
+We generate equally-spaced samples $\boldsymbol{\gamma}_{i}\in\mathbb{R}^{D_{\boldsymbol{\gamma}}}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to 00 and then to a particular $\boldsymbol{\gamma}_{b}$.
+We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center.
+For visual comparison, we show an analogous interpolation from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$ by using a linear geometric interpolation in junction space ${\cal G}$, the expressions for which are in the supplement.
+Fig.
+S4: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$.
+Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$.
+The embedding has learned to be smooth and have an intuitive zero.
+S4.3  Evolution of the Outputs
+Figure S5 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement.
+Specifically, we visualize the result of slicing $b_{k}(x;\mathbf{g})$, which we then apply a non-linearity to in order to amplify the less prominent boundaries.
+We see that early iterations are exploratory and unstructured, and that later iterations reach agreement.
+Fig.
+S5: Evolution of boundaries during iterations, in reading order.
+Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
+S4.4  Qualitative Results for Natural Images
+Fig.
+S6: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [19] and Pidinet [23] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [3], and Field of Junctions [25] with patch size $11$.
+In Figures S7 and S6, we show how the model behaves on noiseless natural images that contain texture and recognizable objects.
+In particular, Figure S6 emphasizes how the boundary maps produced by our model qualitatively differ from those of many classical bottom-up edge-detectors and also from those of learned, end-to-end models that have been trained to match human annotations.
 The figure compares our output to that from Canny [3], Field of Junctions [25] with a patch size of $11$, Pidinet [23], and EDTER [19].
@@
 It does not include mechanisms for grouping based on local texture statistics, nor based on non-local shape and appearance patterns that have semantic meaning to humans.
-Compared to the bottom-up methods of Canny [4] and Field of Junctions [33], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of field $\mathbf{p}[n]$.
+Compared to the bottom-up methods of Canny [3] and Field of Junctions [25], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of windowing field $\mathbf{p}[k]$.
 In contrast, the Field of Junctions and Canny both operate at a single pre-determined choice of local size, so they tend to oversegment some places while undersegmenting others.
-Figure 13: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [28] and Pidinet [31] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [4], and Field of Junctions [33] with patch size $11$.
-Figure 14: Qualitative behavior of our model on noiseless natural images.
-From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
-5 . Additional Examples for Low-light Images
-Figure 15 shows examples of applying our model to indoor images taken by an iPhone XS in low light conditions.
-Figure 15: Visualization of our model’s output for low-light images captured by an iPhone XS.
-From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
-Figure 16 provides additional comparisons for a sample of varying-noise images from the ELD dataset [34].
+Fig.
+S7: Qualitative behavior of our model on noiseless natural images.
+From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}[n]$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
+S5 Additional Examples for Low-light Images
+Figure S8 shows examples of applying our model to indoor images taken by an iPhone XS in low light conditions.
+Figure S9 provides additional comparisons for a sample of varying-noise images from the ELD dataset [26].
 When detecting boundaries at low signal-to-noise ratios, it is difficult to accurately discern finer structures as the noise level increases.
@@
 Our model reduces the severity of this trade-off by automatically adapting its local windowing functions in ways that have learned to account for both the amount of noise and the local geometry of the underlying boundaries.
-In Figure 16 we see that our model is able to capture the double-contour shape of the curved, thin black bars, and that it continues to resolve them as the noise level increases, more than the other low-level methods.
-We also note that only the low-level models resolve this level of detail in the first place: The models trained on human annotations—EDTER, HED, Pidinet, and Structured Forests—miss the double contour entirely, estimating instead a single thick curve.
+In Figure S9 we see that our model is able to capture the double-contour shape of the curved, thin black bars, and that it continues to resolve them as the noise level increases, more than the other low-level methods.
+We also note that only the low-level models resolve this level of detail in the first place: The models trained on human annotations—UAED [28], EDTER, HED, Pidinet, and Structured Forests—miss the double contour entirely, estimating instead a single thick curve.
 We emphasize again that a user can adjust the behavior of Canny and Field of Junctions by tuning their local size parameters, either the filter size for Canny or the patchsize for Field of Junctions.
@@
 Neither system provides the ability to estimate fine grained details and withstand noise, like our model does.
-Figure 16: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise.
-We compare to end-to-end models that are trained to match human annotations (EDTER [28], HED [35], Pidinet [31], and Structured Forests [10]) in addition to low-level models that are not (Canny [4], and Field of Junctions (FOJ) [33]).
-Figure 20 contains additional examples of images cropped from the ELD dataset. Here we include examples with even higher levels of noise to show the complete degradation of our algorithm and others.
-Figure 17:
-Figure 18: (cont.)
-Figure 19: (cont.)
-Figure 20: (cont.) Additional qualitative comparisons between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, using crops from the ELD dataset under increasing amounts of photographic noise, including very high levels of noise.
-6 . Additional Uses of Our Model
+Figure S10 contains additional examples of images cropped from the ELD dataset. Here we include examples with even higher levels of noise to show the complete degradation of our algorithm and others.
+Fig.
+S8: Visualization of our model’s output for low-light images captured by an iPhone XS.
+From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
+Fig.
+S9: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise.
+We compare to end-to-end models that are trained to match human annotations (UAED [28], EDTER [19], HED [27], Pidinet [23], and Structured Edges (SE) [7]) in addition to low-level models that are not (Canny [3], gPb [14], and Field of Junctions (FoJ-17) [25]) with patch size 17.
+Fig. S10:
+Fig. S10: (cont.)
+Fig. S10: (cont.)
+Fig.
+S10: (cont.) Additional qualitative comparisons between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, using crops from the ELD dataset under increasing amounts of photographic noise, including very high levels of noise.
+S6 Additional Uses of Our Model
 Here we demonstrate to uses of our model that follow directly from its output: hole-filling in RGBD images and non-photorealistic stylization.
-S6.1 . Color-based Depth Completion
-Figure 21 shows an example of using our model for simple hole-filling in the depth channels of RGBD images from the Middlebury Stereo Datasets [30, 29].
+S6.1  Color-based Depth Completion
+Figure S11 shows an example of using our model for simple hole-filling in the depth channels of RGBD images from the Middlebury Stereo Datasets [20, 21].
 We run our model on the RGB channels, and then for each pixel $n$ that has a missing depth value, we use our model’s output local attention kernels $a_{n}(x)$ to fill in that pixel’s value using an attention-weighted average of the observed depth values around it.
 This simple algorithm can be applied whenever the hole sizes are smaller than the maximum diameter of our attention maps, which is $34\times 34$ in our current implementation).
-Figure 21: Using our model for depth completion in RGBD images.
+Fig.
+S11: Using our model for depth completion in RGBD images.
 Left: Input RGB channels.
@@
 Right: Completed depth using our model’s output attention kernels.
-S6.2 . Application: Photo Stylization
-Figure 22 shows examples of using our model’s output for image stylization, by superimposing an inverted copy of the output boundary map $\bar{b}_{\eta}[n]$ onto the smoothed colors $\bar{\mathbf{f}}[n]$.
-Figure 22: Examples of stylized natural photographs, created by imposing our method’s output boundary map onto the output smoothed colors.
+S6.2  Application: Photo Stylization
+Figure S12 shows examples of using our model’s output for image stylization, by superimposing an inverted copy of the output boundary map $\bar{b}_{\eta}[n]$ onto the smoothed colors $\bar{\mathbf{f}}[n]$.
+Fig. S12: Examples of stylized natural photographs, created by imposing our method’s output boundary map onto the output smoothed colors.

### v3

@@
-Boundary Attention: Learning to Localize Boundaries under High Noise
+Boundary Attention: Learning curves, corners, junctions and grouping
 Abstract
-We present a differentiable model that infers explicit boundaries, including curves, corners and junctions, using a mechanism that we call boundary attention.
-Boundary attention is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specify an unrasterized description of the local boundary structure in every overlapping patch within an image.
-It operates in a bottom-up fashion, similar to classical methods for sub-pixel edge localization and edge-linking, but with a higher-dimensional description of local boundary structure, a notion of spatial consistency that is learned instead of designed, and a sequence of operations that is end-to-end differentiable.
-We train our model using simple synthetic data and then evaluate it using photographs that were captured under low-light conditions with variable amounts of noise.
-We find that our method generalizes to natural images corrupted by real sensor noise, and predicts consistent boundaries under increasingly noisy conditions where other state-of-the-art methods fail.
+We present a lightweight network that infers grouping and boundaries, including curves, corners and junctions.
+It operates in a bottom-up fashion, analogous to classical methods for sub-pixel edge localization and edge-linking, but with a higher-dimensional representation of local boundary structure, and notions of local scale and spatial consistency that are learned instead of designed.
+Our network uses a mechanism that we call boundary attention: a geometry-aware local attention operation that, when applied densely and repeatedly, progressively refines a pixel-resolution field of variables that specify the boundary structure in every overlapping patch within an image.
+Unlike many edge detectors that produce rasterized binary edge maps, our model provides a rich, unrasterized representation of the geometric structure in every local region.
+We find that its intentional geometric bias allows it to be trained on simple synthetic shapes and then generalize to extracting boundaries from noisy low-light photographs.
 1 Introduction
@@
 However, the inverse problem in computer vision of robustly inferring precise, unrasterized contours from discrete images remains an open challenge, especially in the presence of noise.
-Earlier work by Canny [3] and many others [5, 16, 10, 9, 15] explored the detection of parametric edge models, and there are a variety of bottom-up algorithms that try to connect them by encouraging geometric consistency via edge-linking or message-passing.
-But since the dawn of deep learning, boundaries have almost exclusively been represented using discrete, rasterized maps, and spatial-consistency mechanisms that were previously based on explicit curve geometry have largely been replaced by black-box neural networks.
+Earlier work by Canny [2] and many others [4, 15, 9, 8, 14] explored the detection of unrasterized parametric edge models, and there are a variety of bottom-up algorithms that try to connect them by encouraging geometric consistency via edge-linking or message-passing.
+But since the dawn of deep learning, boundaries have almost exclusively been represented using discrete, rasterized maps; and spatial-consistency mechanisms that were previously based on explicit curve geometry have largely been replaced by black-box neural networks.
 In this paper, we take inspiration from early computer vision work and revisit the task of finding unrasterized boundaries via bottom-up geometric processing.
-But unlike early work, we want to leverage the speed and power of neural networks to learn these processes instead of hand-crafting them.
-We focus our attention on the low-level task of finding boundaries that separate regions of uniform color, as depicted by the toy problem in Figure 1.
-This task becomes difficult at high noise levels, especially within local patches, and we address it by creating a model that can learn to exploit a wide range of low-level cues, such as curves being predominantly smooth with low-curvature; curves tending to meet at corner and junction points; contours tending to have consistent contrast polarity throughout their extent; colors tending to vary smoothly at locations away from boundaries; and so on.
+But unlike early work, we leverage self-attention to learn these processes instead of hand-crafting them, thereby combining the benefits of geometric modeling with the efficiency and representational power of deep learning.
+We focus on the low-level task of finding boundaries that separate regions of uniform color, as depicted by the toy problem in Figure 1.
+This task becomes difficult at high noise levels, especially within local patches, and we address it by creating a model that can learn to exploit a wide range of low-level cues, such as curves being predominantly smooth with low-curvature; curves tending to meet at corner and junction points; contours tending to have consistent contrast polarity throughout their extent; and colors tending to vary smoothly at locations away from boundaries.
 The core of our model is a mechanism we call boundary attention.
@@
 This gives our model enough capacity to learn to localize boundaries with high precision, and to regularize them in the presence of noise without erroneously rounding their corners or missing their junction points.
-Figure 1: Process overview.
+Figure 1: Pipeline overview.
 The image unfolds into stride-1 patches, and boundary attention operates iteratively on their embeddings to produce for each patch: (i) a parametric three-way partitioning, and (ii) a parametric windowing function that defines its effective patch size.
 (Figure 2 shows parameterization details.) This output field implies a variety of global maps, shown in clockwise order: a boundary-aware smoothing of the input colors; an unsigned boundary-distance map; a boundary map; and a map of spatial affinities between any query point and its neighbors.
-We design our model so that all of its components are local and invariant to discrete spatial shifts, enabling it to be trained on small-sized images and then deployed on much larger and differently-shaped ones.
-By tying the weights across multiple portions of the network, we increase its resilience to noise which we validate with ablations.
-Our resulting model is very compact, comprising only 207k parameters, and it runs several times faster than many alternatives.
+We intentionally design our model with components that are local and invariant to discrete spatial shifts, enabling it to be trained on small-sized images and then deployed on much larger and differently-shaped ones.
+By tying the weights across multiple portions of the network, we increase its resilience to noise, which we validate through ablations.
+Our resulting model is very compact, comprising only 207k parameters, and it runs much faster than comparable optimization-based approaches.
 Further, our model can be trained to a useful state with very simple synthetic data, made up of random circles and triangles that are uniformly colored and then corrupted by noise.
-Despite this simplicity, we find that our model can generalize to captured photographs and novel noise processes surprisingly well.
-We evaluate our model on real noisy images and show that it can capture fine details that other methods miss, and moreover maintain its predictions over high levels of noise.
+Despite the simplicity of this training data, we find that the model’s learned internal activations exhibit intuitive behavior, and that the model generalizes to real world photographs, including those taken in low-light and have substantial shot noise.
 Our main contributions can be summarized as follows:
   1. 1.
-    We propose a novel network design that explicitly models boundaries using a boundary attention mechanism that can be used in any deep learning framework.
+    We introduce a bottom-up, feedforward network that decomposes an image into an underlying field of local geometric primitives that explicitly identify curves, corners, junctions and local grouping.
   2. 2.
-    We demonstrate that our model based on boundary attention is more effective than current state-of-the-art methods at precisely and consistently finding boundaries at high levels of sensor noise (Figure 6).
+    We do this by introducing a new parameterization for local geometric primitives, and a new self-attention mechanism that we call boundary attention.
+  3. 3.
+    We identify the best architecture among a family of candidates, and we show that it generalizes from simple synthetic images to real low-light photographs.
 Figure 2: Parameterization details.
@@
 The windowing parameters $\mathbf{p}=(p_{1},p_{2},p_{3})$ are convex weights over a dictionary of binary pillboxes.
-See the supplement for an interactive visualization.
 2 Related Work
@@
 Our model is strongly inspired by this work, and it improves upon it by introducing an enhanced family of parameterized geometric primitives (Figure 2) and by replacing its hand-designed objective and associated non-convex optimization routine with a learned model that is fast, feed-forward and differentiable.
-We also tackle the challenge of representing images that vary in scale across different areas.
-Unlike many low-level methods that assess image edges at a single, frequently manually chosen scale, our approach considers variations in scale throughout the image.
-The field of junctions, for example, requires setting a single patch size that junctions are optimized over, where larger patch sizes allow for better resilience to noise at the expense of memory and detail preservation.
-In contrast we equip each local region with a function to modulate its spatial extent.
-Our work is very different from the recent trend in computer vision of using deep networks to detect boundaries that are semantically meaningful. Examples include HED [27], EDTER [19], and UAED [28].
-These models are trained using annotated segmentation datasets like BSDS500 [2] and NYUDv2 [22], and they aim to identify boundaries between semantic objects (e.g., people, animals) while ignoring boundaries that are inside of these objects (e.g., stripes on a shirt).
-Their large capacities and their end-to-end training allow them to exploit not only low-level cues but also high-level cues such as object and scene familiarity.
-In contrast to this recent trend, we follow earlier work by focusing entirely on identifying boundaries from the bottom up, using low-level cues alone.
-Since our model is trained using simple synthetic data, it does not exploit object familiarity or learn to ignore intra-object edges.
-We find that this approach maintains some advantages: instead of producing a pixelized boundary map, our model produces a rich field of unrasterized primitives that provide better precision as well as information about local grouping.
-Moreover, despite not emphasizing top-down information, we find it performs better than deep models in the presence of high noise levels.
-This suggests that future work may benefit from incorporating some of our architectural elements into larger and higher-capacity systems, in particular to improve interpretability, efficiency, generalization and task-transfer.
+We also tackle the challenge of representing images that vary in scale across different areas: whereas FoJ requires setting a single patch size for the entire image, our model learns to choose patch sizes adaptively in response to the input image.
+We emphasize that our work is very different from a recent trend in computer vision of using deep networks to detect boundaries that are semantically meaningful (e.g.[21, 18, 27]).
+These models aim to identify boundaries between semantic concepts, such as people and animals, while ignoring boundaries that are inside of these concepts, such as stripes on a shirt.
+In contrast to this trend, we follow earlier work by focusing entirely on identifying boundaries from the bottom up, using low-level cues alone.
+Since our model is trained using simple synthetic data, it does not exploit object familiarity or learn to ignore intra-object contrast.
+This approach has advantages: It is not specialized to any predetermined set of semantic concepts or tasks, and instead of producing a pixelized boundary map, it produces a field of unrasterized primitives that provide better precision as well as richer information about local grouping.
+We leave for future work the exploration of how to combine our bottom-up model with other cues that are top-down and semantic.
 3 Representation
 Our system is depicted in Figure 1.
-It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with stride-1 patch embeddings.
-Critically, each embedding explicitly encodes a tuple of values that specifies the geometric structure and spatial extent of the local boundaries within a patch.
-Because the embeddings are forced through this geometric bottleneck, we refer to our model’s core mechanism as boundary attention.
-We instantiate our bottleneck using a learned linear mapping from the embedding dimension to a predefined lower-dimensional space of unrasterized boundary patterns that we call junction space.
-Junction space has the benefit of specifying local boundary patterns without rasterization and thus with unlimited spatial precision.
-As depicted in Figure 2 and described in [25], it also has the benefit of including a large family of local boundary patterns, including uniformity (i.e., absence of boundaries), edges, bars, corners, T-junctions and Y-junctions.
-We also modulate the spatial extent of each junction by learning an associated windowing function, described in Section 3.1.
-To enable communication between patch embeddings, each of which corresponds to a local region of an image, we leverage the idea that each region of the image is described by multiple patches, and that these patches should agree on compatible representations of the global image structure.
-The ideal embedding for each patch should aptly represent its local geometry while agreeing with the structural claims of neighboring patches.
-The next section introduces our parameterization of junction space and some associated operators.
-Then Section 4 describes the architecture that we use to transform an image into its junction space representation.
+It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with $D$-dimensional, stride-1 embeddings.
+Critically, each $D$-dimensional embedding explicitly encodes a tuple of values that specifies the geometric structure and spatial extent of the unrasterized local boundaries within a patch.
+Our model relies on a learned (linear) mapping from the embedding dimension $D$ to a hand-crafted, lower-dimensional space of unrasterized boundary patterns that we call junction space.
+Junction space has the benefit of specifying per-patch boundary patterns without rasterization and thus with unlimited spatial precision.
+As depicted in Figure 2 and described in [23], it also has the benefit of including a large family of per-patch boundary patterns, including uniformity (i.e., absence of boundaries), edges, bars, corners, T-junctions and Y-junctions.
+Different from [23], we additionally modulate the spatial extent of each patch by learning an associated windowing function, to be described in Section 3.1.
+To enable communication between patch embeddings, each of which corresponds to a local patch, we leverage the idea that each image point is covered by multiple patches, and that overlapping patches must agree in their regions of overlap.
+This has the effect of tying neighboring patches together, analogous to cliques in a Markov random field.
+We refer to the core mechanism of our model as boundary attention.
+We introduce it by first defining our hand-crafted parameterization of junction space and some associated operators.
+Then Section 4 describes the architecture that we use to transform an image into a coherent junction-space representation.
 3.1 Boundary Primitives
-Throughout this paper we use parentheses $(x)$ for continuous signals defined on the 2D image plane $[0,W]\times[0,H]$ and square brackets $[n]$ for discrete signals defined on the pixel grid.
-We use $c[n]$ for the coordinates of the $n$th pixel’s center.
-Denote the $K$-channel input image by $\{\mathbf{f}[n]\}$, where $\mathbf{f}[n]\in\mathbb{R}^{K}$ is the vector image value at the discrete pixel grid index $n$.
-Our approach is to analyze the image into a field of dense, stride-1 overlapping local patches, each having a square support $\Omega_{k}(x)$ for the $k$th patch centered at the $k$th pixel.
-There are many ways to partition a local region $\Omega_{k}(x)$, and one can define parametric families of partitions.
+We use parentheses $(x)$ for continuous signals defined on the 2D image plane $[0,W]\times[0,H]$ and square brackets $[n]$ for discrete signals defined on the pixel grid.
+We use $c[n]$ for the coordinates of the pixel with integer index $n$.
+Denote the $C$-channel input image by $\{\mathbf{f}[n]\}$, where $\mathbf{f}[n]\in\mathbb{Q}^{C}$ is the vector image value at the discrete pixel grid index $n$.
+Our approach is to treat the image as a field of dense, stride-1 overlapping local patches.
+We use $\Omega_{k}(x)$ to denote the spatial support of the patch that is centered at the pixel whose integer index is $k$.
+There are many ways to partition a local patch $\Omega_{k}(x)$, and one can define parametric families of partitions.
 For example the set of oriented lines provides a two-parameter family of partitions, with each member of the family separating the region into points that lie on one side of a line or the other.
@@
 We use the notation $\mathbf{g}=(\boldsymbol{u},\theta,\boldsymbol{\omega})$, where $\boldsymbol{u}=(u,v)\in\mathbb{R}^{2}$ is the vertex, $\theta\in\mathbb{S}^{1}$ is the orientation, and $\boldsymbol{\omega}=(\omega_{1},\omega_{2},\omega_{3})$ are barycentric coordinates (defined up to scale) for the three relative angles, ordered clockwise starting from $\theta$.
-Our convention is to express the vertex coordinates relative to the center of region $\Omega_{k}(x)$, which is located at $c[k]$, and we denote that the vertex is free to move outside of this region.
+Our convention is to express the vertex coordinates relative to the center of region $\Omega_{k}(x)$, which is located at $c[k]$, and we note that the vertex is free to move outside of this region.
 We also note that up to two angles $\omega_{j}$ can be zero.
-This all makes it possible to represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions), and do so with unlimited spatial resolution.
+This all makes it possible to smoothly represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions), and do so with unlimited spatial resolution.
 Fixing a value for $\mathbf{g}$ induces three wedge support functions, denoted by
@@
 Analytic expressions for them are included in the supplement.
-In order to enable the size of each region $\Omega_{n}$ to be able to adapt to the local geometry and noise conditions, we equip each one with a parameterized local windowing function $w_{k}(x;\mathbf{p})\in[0,1]$, with parameters $\mathbf{p}\in{\cal P}=\triangle^{W-1}$ that are the coefficients of a convex combination of $W$ square pillbox functions.
+In order to enable the size of each region $\Omega_{k}$ to adapt to the local geometry and noise conditions, we equip each one with a parameterized local windowing function $w_{k}(x;\mathbf{p})\in[0,1]$, with parameters $\mathbf{p}\in{\cal P}=\triangle^{W-1}$ that are the coefficients of a convex combination of $W$ square pillbox functions.
 That is,
 $w_{k}(x;\mathbf{p})=\sum_{i=1}^{W}p_{i}\mathbf{1}[\|x-c[k]\|_{\infty}\leq D_{i}],$ (3)
-[…]
+where $\|\cdot\|_{\infty}$ is the $\ell^{\infty}$-norm, and $\mathbf{1}[\cdot]$ is the indicator function that returns $1$ if the argument is true; and 00 otherwise.
+In our experiments we use $W=3$ and diameters $\mathbf{D}=(3,9,17)$.
+Figure 3 shows some examples.
 Figure 3: Example of our model’s output, with examples from two different regions.
@@
 Bottom row: Four types of global maps that are implied by accumulating values from the output field and rendered patches.
-See the supplement for interactive visualizations.
 3.2 Gather and Slice Operators
@@
 $\bar{\mathbf{f}}[n]$ $=\frac{\sum_{k}w_{k}(c[n];\mathbf{p}[k])\sum_{j}\mathbf{f}_{kj}s_{kj}(c[n];\mathbf{g}[k])}{\sum_{k}w_{k}(c[n];\mathbf{p}[k])\sum_{j}s_{kj}(c[n];\mathbf{g}[k])}.$ (6)
-Note that only patches that contain $c[n]$ contribute to the sum, i.e., the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$.
-The expressions for the pixel-wise distance map variance $\nu_{d}[n]$, and for the feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are defined accordingly and included in the supplement.
-Intuitively, slicing $\mathbf{f}_{kj}$ can be thought of globally accumulating information from each patch’s junction representation.
+Here, the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$ so that only patches that contain $c[n]$ contribute to the sum.
+Similar expressions for pixel-wise distance map variance $\nu_{d}[n]$ and feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are included in the supplement.
+Intuitively, slicing represents an accumulation of regional information into a pixel, as dictated by the partitions of all of the patches that contain the pixel.
 3.3 Visualizing Output
-At our network’s output, we are provided with a field of tuples representing the junction and window parameters $\{(\mathbf{g}[k],\mathbf{p}[k]\}$ for all stride-1 patches of the input image.
-Though these outputs are defined continuously, we can visualize them in Figure 3 by rendering the windowing functions $w_{k}(x;\mathbf{p})$ (second column), and the binary junction wedge supports $s_{kj}(x;\mathbf{g})$ colored purple, orange, and blue (third column), along with the input image patches (first column).
-To the left, we show the input image and the regions from which the patches were extracted from.
-Additionally, we can use the learned junction parameters to render unsigned distance patches $d_{k}(x;\mathbf{g})$ (fourth column), boundary patches $b_{k}(x;\mathbf{g})$ (fifth column), and the binary junction wedges $s_{kj}(x;\mathbf{g})$ colored with their respective gathered colors (sixth column).
-All of the patches are defined continuously, so we can render them at any resolution.
+Our network’s output is a field of tuples representing the junction and windowing parameters $\{(\mathbf{g}[k],\mathbf{p}[k]\}$ for all stride-1 patches of the input image.
+We visualize them in Figure 3 by rasterizing the continuous windowing functions $w_{k}(x;\mathbf{p})$ (second column) and binary wedge supports $s_{kj}(x;\mathbf{g})$, which are colored purple, orange, and blue (third column).
+To the left, we show the input image and the regions from which the patches were extracted.
+Additionally, we can use the output junction parameters to rasterize unsigned distance patches $d_{k}(x;\mathbf{g})$ (fourth column), boundary patches $b_{k}(x;\mathbf{g})$ (fifth column), and wedge supports $s_{kj}(x;\mathbf{g})$ that are re-colored with their respective wedge features $\mathbf{f}_{kj}$ (sixth column).
+Note that all of these are defined continuously, so can be rasterized at any resolution.
 We expect the shapes of junction boundaries in overlapping regions to agree, so that the variances $\nu_{d}[n],\nu_{f}[n]$ are small at every pixel.
@@
 setting $\eta=0.3$.
-The output contains much more information than these fields of means.
-For any query pixel $n$, we can probe the wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ that contain it, thereby obtaining a spatial attention map $a_{k}(x)$ that surrounds the query pixel.
-This is the boundary-aware spatial kernel that turns a neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the value $\bar{\mathbf{f}}[n]$:
+For any query pixel $k$, we can also probe the containing wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ to compute a spatial affinity map $a_{k}(x)$ that surrounds the query pixel.
+This represents the affinity between point $c[k]$ and a surrounding neighborhood with diameter that is twice that of $\Omega(x)$.
+It is also the boundary-aware spatial kernel that turns the neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the gathered value $\bar{\mathbf{f}}[n]$ via
 $\bar{\mathbf{f}}[n]=\sum_{k}a_{n}(c[k])\mathbf{f}[k].$ (8)
-The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6, and its maximum diameter is twice that of $\Omega(x)$.
+The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6.
 The spatial affinity maps for two probed points are shown in the leftmost image of the bottom row of Figure 3.
-Like slicing, probing the affinity map around a pixel is a form of global accumulation that takes values defined per-patch and synthesizes them into a global representation.
+Like slicing, querying the affinity map at a pixel is a form of regional accumulation from patches to a pixel.
 Figure 4: Model Architecture.
@@
 The first iteration uses $\boldsymbol{\gamma}^{0}[n]=\boldsymbol{\gamma}_{0}[n]$, $\bar{\mathbf{f}}^{0}[n]=\mathbf{f}[n]$, and $\boldsymbol{\pi}^{0}[n]=\boldsymbol{\pi}_{o}$ with $\boldsymbol{\pi}_{o}$ learned across the training set.
-Boundary attention repeats $T=8$ times, with one set of weights for the first four iterations and another for the last four iterations.
-There are $207$k trainable parameters in total.
+Boundary attention repeats $T=8$ times, with one set of weights for the first four iterations and a separate set of weights for the last four iterations, resulting in 207k trainable parameters total.
 4 Network Architecture
-We seek a differential architecture that can effectively initialize and then refine the fields $\{(\mathbf{g}^{t}[k],\mathbf{p}^{t}[k])\}$ using a manageable number of iterations.
-This bears similarities to the edge localization and edge-linking steps of early edge detectors like Canny’s, but with a more sophisticated local boundary model, and with mechanisms for spatial consistency that are learned instead of designed.
+We seek a differentiable architecture that can effectively initialize and then refine the fields $\{(\mathbf{g}^{t}[k],\mathbf{p}^{t}[k])\}$ using a manageable number of iterations.
+This approach is motivated by the edge localization and edge-linking steps of early edge detectors like Canny’s, but has a more sophisticated local boundary model, and mechanisms for spatial consistency that are learned instead of designed.
 It is also analogous to the original field of junctions algorithm [23], which uses coordinate descent for initialization and iterations of gradient descent for refinement.
@@
 This means that our model can be trained on small images and then deployed on much larger ones.
-Also, our model has only $207$k learnable parameters, making it orders of magnitude smaller than most deep semantic boundary detectors.
-As a point of reference, the full EDTER model has 109M parameters [19].
+Also, our model has only $207,000$ learnable parameters, making it orders of magnitude smaller than most deep semantic boundary detectors.
+As a point of reference, Diffusion Edge, a recent diffusion-based semantic boundary detection model uses on the order of 300 million parameters [26].
 Our model is depicted in Figure 4.
-It implicitly uses higher-dimensional embeddings of the field elements, $\boldsymbol{\gamma}^{t}[k]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[k]\in\mathbb{R}^{D_{\pi}}$, which can be decoded at any iteration using the learned linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$.
-Our final model uses $D_{\gamma}=64$ and $D_{\pi}=8$ which, as shown in the supplement, provides enough capacity to learn a smooth latent representation of junction space.
-Use separate embeddings for the junction and windowing fields provides a disentangled representation of both.
+It represents the field elements as higher-dimensional embeddings $\boldsymbol{\gamma}^{t}[k]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[k]\in\mathbb{R}^{D_{\pi}}$, which can be decoded at any iteration using the learned linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$.
+Our final model uses $D_{\gamma}=64$ and $D_{\pi}=8$ which we show in our experiments provides enough capacity to learn a smooth latent representation of junction space.
+Using separate embeddings for the junction and windowing fields provides a disentangled representation of both.
 Given an input image, the network first applies a “neighborhood MLP-mixer”, which is a modified MLP-Mixer [22] that replaces the global spatial operations with convolutions of kernel size $3$.
@@
 This hidden state is then refined by a sequence of eight boundary attention iterations, which we describe next.
-(See the supplement for a visualization of the decoded hidden states as they evolve.)
-The eight iterations of refinement are broken into two boundary attention blocks, each with learned weights.
-To process our input, we first add in a linear projection of the initial hidden state.
-This is essentially a skip connection that allows our network to retain information from the initial hidden state estimate at later stages of processing.
-Next, we copy our hidden state into two identical pieces.
-We concatenate a dimension 8 learned windowing embedding to one of the copies and the input image plus the current estimate of the smoothed global features to the other.
-We then do neighborhood cross-attention: each pixel in the first copy does two iterations of cross attention with a size 11 patch of the second copy.
+(See our experiments for a visualization of the decoded hidden states as they evolve.)
+The eight iterations of refinement are broken into two blocks, each with learned weights.
+In each iteration, we first add a linear mapping of the initial hidden state to the current hidden state, which acts as a skip connection.
+Next, we clone our hidden state, concatenating a dimension 8 learned windowing embedding to one of the copies and the input image plus the current estimate of the smoothed global features to the other.
+We treat the copies as the inputs to neighborhood cross-attention: each pixel in the first copy does two iterations of cross attention with a size 11 patch of the second copy.
 We add a learned $11\times 1$ 11 positional encoding to the patch, which allows our network to access relative positioning, even though global position cues are absent.
 We follow each self attention layer with a small MLP.
-To transform our output or intermediary hidden state into junction space and render output images, we use a simple linear mapping.
+To transform our output or intermediary hidden state into junction space and render patch visualizations, we use a simple linear mapping.
 We separate the windowing embedding (the last 8 dimensions) from the junction embedding (the first 64 dimensions) and project each through a linear layer.
-We map the hidden state to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$.
+We map the junction embedding to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$.
 These serve as the inputs to our gather and slice operators.
@@
 4.1 Training
-We train our network in two stages: we begin by training the neighborhood MLP-mixer with the first Boundary Attention Block, and then we add the second block and retrain end-to-end.
-Because junction space is non-convex and difficult to optimize over, this two step training procedure helps the network converge.
-We apply our loss to the two final iterations of our network, so for the first stage of training we apply a loss to iterations 3 and 4, and for end-to-end refinement we apply a loss to iterations 7 and 8.
-We weigh the final loss three times higher than the second to last loss, which encourages the network to allocate capacity to producing high quality outputs, while still providing supervision to intermediate junction estimates.
-We train our method using a combination of four global losses applied to global (i.e.
+Estimating the best junctions for noisy image patches is a non-convex optimization that is prone to getting stuck in local minima [23].
+We find that training our network in three stages with increasingly complex synthetic training data produces the best model.
+First we train the network on small $21\times 21$ images containing a single junction corrupted by low amounts of Gaussian noise.
+Upon convergence, we retrain the network on larger $100\times 100$ images containing a single circle and triangle corrupted by moderate Gaussian noise.
+Finally, we retrain the network on a high-noise synthetic dataset containing $125\times 125$ images consisting of many overlapping triangles and circles.
+(See supplement for examples.)
+We calculate our loss using the outputs of the two final iterations of our network, where the loss of the final output is weighted three times as much as the loss applied to the output before it.
+This encourages the network to allocate capacity to producing high quality outputs, while still providing supervision to intermediate junction estimates.
+We train our model using a combination of four global losses applied to global (i.e.
 sliced) fields, and two patch-wise losses applied to individual patches.
-The first two losses are supervision losses penalizing mismatches between our network’s predictions and the ground truth feature and boundary maps:
+The first two losses are supervision losses penalizing mismatches between our network’s predictions and the ground truth feature and distance maps:
 $\mathcal{L}_{f}$ $=\sum_{n}\alpha[n]\|\bar{\mathbf{f}}[n]-\mathbf{f}_{\text{GT}}[n]\|^{2},$ (9)
 $\mathcal{L}_{d}$ $=\sum_{n}\alpha[n]\left(\bar{d}[n]-d_{\text{GT}}[n]\right)^{2},$ (10)
-where $\mathbf{f}_{\text{GT}}$ and $d_{\text{GT}}$ are the ground truth features and distance maps, respectively, and $\alpha[n]$ is a pixel importance function defined as:
-$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+(1-C),$ (11)
+where $\mathbf{f}_{\text{GT}}$ and $d_{\text{GT}}$ are the ground truth features and distance maps, respectively, and $\alpha[n]$ is a pixel importance function defined as
+$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+C,$ (11)
 with $\beta$ and $C$ controlling how much weight to give pixels near boundaries.
-We set $\beta=0.1$, $\delta=1$, and anneal $C$ to give more weight to non-boundary locations later in training.
-We also tested a more involved pixel importance mask that gave extra weight to visible vertices and intersections, the details for which can be found in the supplement.
-Note that using noiseless feature maps for supervision in Equation 9 encourages the windowing functions to be large in smooth regions.
-On top of the two supervision losses we apply two consistency losses from [25], that minimize the per-pixel variances $\nu_{f}[n]$ and $\nu_{d}[n]$.
-Similar to the supervision losses, we weigh those by $\alpha$ from Equation 11.
+We set $\beta=0.1$, $\delta=1$, and $C=0.3$.
+(The supplement contains additional tests with a more involved importance mask.) Using noiseless feature maps for supervision in Equation 9 has the effect of encouraging windowing functions to be as large as possible, because larger regions imply averaging over a greater number of noisy input pixel values.
+On top of the two supervision losses we apply two consistency losses borrowed from [23], which minimize the per-pixel variances $\nu_{f}[n]$ and $\nu_{b}[n]$.
+Here, we weight them by $\alpha[n]$ from Equation 11.
 These consistency losses encourage the junction shapes $\mathbf{g}$ in overlapping regions to agree.
-Minimizing $\nu_{f}[n]$ also encourages windowing functions to be large, because that increases the gather area, which in turn reduces noise in wedge features $\mathbf{f}_{nj}$ that are sliced to compute the variance $\nu_{f}[n]$ at each $n$.
+Minimizing $\nu_{f}[n]$ also provides a second mechanism to encourage windowing functions to be large.
+Larger windows increase gather area, thereby reducing noise in wedge features $\mathbf{f}_{nj}$ that are sliced to compute variance $\nu_{f}[n]$ at each pixel $n$.
 Finally, we use two patch-wise losses to encourage individual feature and distance patches to agree with the supervisory ones:
@@
 $\chi[k]=\left(\sum_{n\in\Omega_{k}}(d_{\text{GT}}[n]+\delta^{\prime})\right)^{-1},$ (14)
-and we set $\delta^{\prime}=1$. These per-patch losses gives the network a more direct signal on how to adjust its weights than purely global losses, which average over multiple patches.
+with $\delta^{\prime}=1$. These per-patch losses provide a more direct signal for adjusting model weights, compared to per-pixel losses which average over multiple patches.
 Figure 5: Left: ODS F-score for our method and multiple baselines at different noise levels computed on noisy synthetic data.
@@
 Our method has the best average F-score while also being much faster than the second best method Field of Junctions.
-Figure 6: Left: Qualitative comparison of our model versus other methods for noisy crops from the ELD dataset [26].
+Figure 6: Left: Qualitative comparison of our model versus other methods for a noisy crop from the ELD dataset [24].
+See the supplement for more examples.
 Our method provides more detail than other techniques at low noise and is more robust to high levels of noise.
@@
 Implementation details.
-We train our model on noisy synthetic data of colorful combinations of triangles and circles.
+In the final stage of training, we use noisy synthetic data of many randomly-colored overlapping triangles and circles.
 We render $240\times 320$ images containing $15$ to $20$ shapes each, but use $125\times 125$ crops for training.
-To those crops we add Gaussian and/or Perlin noise [17], and with probability $0.1$ we average over the color channels to produce grayscale inputs.
+To those crops we add Gaussian or Perlin noise [16], and with probability $0.1$ we average over the color channels to produce grayscale inputs.
 Our dataset contains $10^{5}$ images, $90\%$ of which are used for training, and the rest for validation.
-For a visualization of training data and training and optimization details, refer to the supplement.
 Baselines.
-Since we focus on low-level edge detection, we primarily compare against other techniques with a similar focus and that also do not train on large semantic datasets: Canny [3], gPb [14], Structured Edges (SE) [7], and Field of Junctions (FoJ) [25].
-However, since most recent work tends to focus on semantic edge detection, in the interest of being comprehensive, we also provide a comparison to several modern works: UAED [28], EDTER [19], and HED [27].
-We note that this is not a clean comparison due to the difference in tasks – predicting edges with low-level cues vs predicting edges with semantic meaning.
-Quantitative results and inference times across different noise levels.
-For quantiative evaluation, our focus on low-level cues means that we cannot use semantic edge detection benchmarks like BSDS50 [2] since our method and other low-level baselines will both predict edges that are not semantically meaningful and fail to predict edges that are semantically meaningful but lack contrast.
-We instead rely on synthetic data where the ground truth edges can be determined with perfect precision and any level of noise can be added.
-In this experiment, we use geometric objects and per-pixel additive Gaussian noise with a random variance.
-Figure S3 compares the performance and inference time of our method and baselines under different noise levels.
+Since we focus on bottom-up edge and corner detection, we compare against other techniques with a similar focus and that, like us, do not train on large semantic datasets: Canny [2], gPb [13], Structured Edges (SE) [6], and Field of Junctions (FoJ) [23].
+Quantitative results.
+For quantitative evaluation we cannot use semantic edge detection benchmarks like BSDS500 [1], because our model and the comparisons are designed to predict all edges, including those that are not semantically meaningful.
+We instead rely on synthetic data, where the ground truth edges can be determined with perfect precision, and inputs can be controllably noised.
+Our evaluation data comprises geometric objects and per-pixel additive Gaussian noise with a random variance.
+Figure 5 compares the performance and inference time of our method and baselines under different noise levels.
 The tuneable parameters for Field of Junctions were chosen to maximize its performance on noisy images with $17\times 17$ patches.
-Notably our method’s adaptive windowing function gives it at edge compared to the Field of Junctions at low noise, enabling it to capture finer details, with only slightly worse performance under extreme noise conditions.
-Our method is also orders of magnitude faster than FoJ as shown in the plot to the right.
+Notably our method’s adaptive windowing function gives it an edge compared to the Field of Junctions at low noise, enabling it to capture finer details, with only slightly worse performance under extreme noise conditions.
+Our method is also orders of magnitude faster than FoJ, as shown at right.
 Qualitative results on real images.
-As shown in Figure 6, despite being trained on synthetic data, our method can outperform existing state-of-the-art methods at multiple levels of real sensor noise present in ELD [26].
+As shown in Figure 6, despite being trained on synthetic data, our method can detect edges in real photographs with multiple levels of real sensor noise present in ELD [24].
 Our method produces crisp and well-defined boundaries despite high levels of noise.
-Figure 7 shows that our method makes reasonable boundary estimates for even complex in-the-wild images.
-Our supplement contains additional examples.
-Figure 7: Our model can find compelling boundaries in natural in-the-wild images, despite training on simple synthetic data.
-Left image from the SIDD [1] dataset; right image taken with a phone camera.
+The supplement includes additional examples that show that our method makes reasonable boundary estimates for other captured images.
 Repeatability on real images.
-We quantify each model’s noise resilience on real low-light images by measuring the repeatability of its boundaries over a collection of scenes from the ELD dataset.
+We quantify each method’s noise resilience on real low-light images by measuring the repeatability of its boundaries over a collection of scenes from the ELD dataset.
 For each scene, we run the model on the lowest-noise image and then measure, via ODS F-score, how much its predictions change with increasing noise level.
 Figure 6 shows the averages of these scores across the collection.
-Our model provides more consistent results for increasingly noisy images than other methods.
-6 Ablations
-[…]
-table:ablation1 shows what happens when the initial per-patch Neighborhood MLP-mixer [NM] is used alone, versus combining it with one boundary attention block [BA-1] or two such blocks [BA-2].
-Tying the attention weights across iterations provides a slight advantage at higher noise levels, with only a slight penalty in accuracy at lower noise.
-Our final model is boxed.
-It provides the best overall performance, accepting slightly lower accuracy at low noise in exchange for better accuracy at higher noise levels.
-[…]
-[…]
+Our model provides more consistent results for increasingly noisy images than other methods, in addition to capturing fine details that other methods miss.
+Properties of learned embedding.
+We find that our model learns a spatially smooth embedding $\boldsymbol{\gamma}\in\mathbb{R}^{D_{\gamma}}$ of junction space $\mathbf{g}\in{\cal G}$.
+In Figure 7 we generate equally-spaced samples $\boldsymbol{\gamma}_{i}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to $\boldsymbol{0}$ and then to a particular $\boldsymbol{\gamma}_{b}$; and then to each sample we apply the learned embedding to compute and visualize the implied junction $\mathbf{g}_{i}$.
+We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center.
+For visual comparison, we show an analogous geometric interpolation in junction space ${\cal G}$ (see the supplement for expressions) from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$.
+Figure 7: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$.
+Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$.
+The embedding has learned to be smooth and have an intuitive zero.
+Figure 8: Evolution of boundaries during iterations, in reading order.
+Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
 […]
@@
 Table 1: Ablations: impact of component combinations and weight-tying on F-score (higher is better). Our reported model is boxed.
+Evolution.
+Figure 8 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement.
+Specifically, we visualize the result of slicing $b_{k}(x;\mathbf{g})$, to which we apply a non-linearity that amplifies less-prominent boundaries.
+We see that early iterations are exploratory and unstructured, and that later iterations agree.
+6 Ablations
+Table 1 shows what happens when the initial per-patch Neighborhood MLP-mixer [NM] is used alone, versus combining it with one boundary attention block [BA-1] or two such blocks [BA-2].
+Tying the attention weights across iterations provides a slight advantage at higher noise levels, with only a slight penalty in accuracy at lower noise.
+Our final model (boxed) provides the best overall performance, accepting slightly lower accuracy at low noise in exchange for better accuracy at higher noise levels.
+Table 2 compares many other variations of the boxed model from Tab.
+1, with asterisks indicating the model specifications used in the paper.
+Iterations varies the number of attention-iterations within each block, and Neighborhood varies the attention neighborhood size.
+MLP Input varies the features concatenated to the hidden state prior to the pre-attention MLP and shows that replacing the gathered colors $\bar{\mathbf{f}}^{t}$ with a constant array or the input image values $\mathbf{f}$ performs worse.
+Windowing shows that using fixed, square windowing functions $w_{n}(x)$ of any size performs equal (at low noise) or worse than inferring them adaptively.
+Somewhat intuitively, a small $9\times 9$ patch size performs well at low noise, but performance lags under noisy conditions where using a larger patch size increases the spatial extent of the communication across patches.
 […]
-Iterations 2 $0.857$ $0.611$ $0.307$
-3 $0.867$ $0.647$ $0.329$
+Iterations 3 $0.867$ $0.647$ $0.329$
 4* 0.872 0.673 0.348
 5 $0.871$ $0.663$ $0.337$
-Neighborhood $3\times 3$ $0.835$ $0.531$ $0.240$
-$7\times 7$ $0.869$ $0.637$ $0.313$
+Neighborhood $7\times 7$ $0.869$ $0.637$ $0.313$
 $9\times 9$ $0.871$ $0.658$ $0.325$
@@
 Avg. features* 0.872 0.673 0.348
-Loss Uniform $0.844$ $0.629$ $0.313$
-Weighted* 0.872 0.673 0.348
-Windowing Fixed - $3\times 3$ $0.834$ $0.546$ $0.205$
-Fixed - $9\times 9$ 0.872 $0.662$ $0.328$
-Fixed - $17\times 17$ $0.859$ $0.662$ $0.328$
+Windowing Fixed 0.872 $0.662$ $0.328$
 Inferred* 0.872 0.673 0.348
 Table 2: Ablations: variations of our reported model.
-We compare the number of iterations of boundary attention, the attention neighborhood size, the features concatenated to the hidden state as input to the MLP, the impact of a spatially weighted loss, and a fixed versus learned windowing function.
+We compare the number of iterations of boundary attention, the attention neighborhood size, the features concatenated to the hidden state as input to the MLP, and a fixed versus learned windowing function.
 Asterisks indicate our reported choices.
@@
 Our bottom-up, feedforward network can encode images of any resolution and aspect ratio into a field of geometric primitives that describe the local structure within every patch.
-Despite being trained on simple geometric synthetic data, our method can consistently localize boundaries under high levels of noise better than state-of-the-art deep networks, while achieving similar performance to recent optimization-based approaches in a fraction of the time.
-Supplemental Material for Boundary Attention: Learning to Localize Boundaries under High Noise
+This work represents a step to the goal of synergizing the benefits of low-level parametric modeling with the efficiency and representational power of deep learning.
+Supplemental Material for Boundary Attention: Learning curves, corners, junctions and grouping
 S1 The space of $M$-junctions
@@
 S1.1  Interpolation
-Another advantage of the present parameterization compared to that of the original [25] is that it is a simply-connected topological space and so allows for defining mechanisms for smoothly interpolating between any two junctions $\mathbf{g}$ and $\mathbf{g}^{\prime}$.
-In our implementation we simply define interpolation variable $t\in[0,1]$ and compute interpolated junctions $\mathbf{g}^{(t)}=\{\boldsymbol{u}^{(t)},\theta^{(t)},\boldsymbol{\omega}^{(t)}\}$ using
-$\boldsymbol{u}^{(t)}$ $=(1-t)\boldsymbol{u}+t\boldsymbol{u}^{\prime}$ (24)
-$\tilde{\boldsymbol{\omega}}^{(t)}$ $=(1-t)\tilde{\boldsymbol{\omega}}+t\tilde{\boldsymbol{\omega}}^{\prime},$ (25)
+Another advantage of the present parameterization compared to that of the original [23] is that it is a simply-connected topological space and so allows for defining mechanisms for smoothly interpolating between any two junctions $\mathbf{g}=\left(\boldsymbol{u},\theta,\boldsymbol{\omega}\right)$ and $\mathbf{g}^{\prime}=\left(\boldsymbol{u}^{\prime},\theta^{\prime},\boldsymbol{\omega}^{\prime}\right)$.
+In our implementation we define interpolation variable $t\in[0,1]$ and compute interpolated junctions $\mathbf{g}(t)=\{\boldsymbol{u}(t),\theta(t),\boldsymbol{\omega}(t)\}$ using a simple combination of linear and spherical linear interpolation:
+$\boldsymbol{u}(t)$ $=(1-t)\boldsymbol{u}+t\boldsymbol{u}^{\prime}$ (24)
+$\tilde{\boldsymbol{\omega}}(t)$ $=(1-t)\tilde{\boldsymbol{\omega}}+t\tilde{\boldsymbol{\omega}}^{\prime},$ (25)
 and
-[…]
-$(p,q)$ $=\operatorname{Slerp}\left((\cos\theta,\sin\theta),(\cos\theta^{\prime},\sin\theta^{\prime}),t\right),$
-where $\operatorname{Slerp}()$ is the 2D geometric spherical linear interpolation operator,
-$\operatorname{Slerp}(\boldsymbol{p},\boldsymbol{p}^{\prime},t)=\frac{\sin\left((1-t)\Delta\theta\right)}{\sin\left(\Delta\theta\right)}\boldsymbol{p}+\frac{\sin\left(t\Delta\theta\right)}{\sin\left(\Delta\theta\right)}\boldsymbol{p}^{\prime},$
-with $\Delta\theta=\arccos(\boldsymbol{p}\cdot\boldsymbol{p}^{\prime})$.
-The bottom row of Figure 9 in the main paper visualizes a set of samples from smooth trajectories in junction space using this mechanism.
+$\theta(t)=\theta+t\Delta\theta,$ (26)
+with
+$\Delta\theta=\begin{cases}\theta^{\prime}-\theta-2\pi,&\text{if }\theta^{\prime}-\theta>\pi\\ \theta^{\prime}-\theta+2\pi,&\text{if }\theta^{\prime}-\theta<-\pi\\ \theta^{\prime}-\theta,&\text{otherwise},\end{cases}$
+assuming $\theta,\theta^{\prime}\in[0,2\pi)$. The bottom row of Figure 7 in the main paper visualizes a set of samples from smooth trajectories in junction space using this mechanism.
 S2 Training Data
@@
 The highly-varying types and amounts of noise also encourages the model to use large window functions $w(x;\mathbf{g})$ when possible, since that reduces noise in the gather operation and reduces variance $\nu_{f}[n]$.
-Our dataset, which we call Kaleidoshapes, will be released upon publication, along with the code for generation, training and evaluation.
+Our dataset, which we call Kaleidoshapes, is available publicly, along with the code for generation, training and evaluation.
 Shapes and colors.
@@
 Here we provide the details of the two blue blocks in the main paper’s Figure 3: Neighborhood MLP-Mixer and Neighborhood Cross-attention.
-We implement our model in JAX, and we will publicly share our model code and its pretrained weights upon publication.
+Our model was implemented with JAX and its code and trained weights are available publicly.
 S3.1  Neighborhood MLP-Mixer
@@
 Inside this block are two transformer layers whose cross-attention components are replaced with neighborhood cross-attention components that are restricted to a spatial neighborhood of pixels.
-We use $11\times 1$ 11 neighborhoods in our implementation.
+We use $11\times 1$ 11 neighborhoods in our implementation, which our ablations showed produced the best performance.
 In each neighborhood containing a query token, we add a learned positional encoding to the key/value tokens which is relative to the neighborhood’s center and is the same for all neighborhoods.
@@
 S3.3  Training Details
-We pretrain the neighborhood MLP-mixer and the first boundary attention block on a simplified variation of our Kaleidoshapes dataset, where each image is $100\times 100$ and contains a single triangle and circle with additive zero-mean Gaussian noise.
-We omit the global losses of Equations $9$ and $10$ during this pretraining phase.
+During the first two stages of training where we train our model on junction images and simplified circle triangle images, we omit the global losses of Equations $9$ and $10$.
 This primes the network to learn meaningful hidden states $\boldsymbol{\gamma}[n]$ and prevents the “collapsing” of junctions, where the boundary-consistency loss (i.e.
@@
 Like in pretraining, we add a spatial importance that prioritizes region containing a corner (i.e., a visible triangle vertex) or a junction (i.e., a visible intersection between the boundaries of any two shapes).
-Our trained weights will be released upon publication.
 S4 Model Behavior
-S4.1  Generalization to Unseen Shapes
-During training, our model observes only circles and triangles of varying sizes and colors.
-We measure how well our model generalizes to unseen shapes by creating a new synthetic dataset containing rectangles, polygons, and ellipses of various sizes, colors, and rotations.
-We evaluate our model against low-level methods as well as methods trained on semantic data, and note the results here closely align with the performance trends on kaleidoshapes (Fig.
-5), suggesting that our bottom-up model can generalize to unseen geometries.
-Fig. S3: We measure how well our model can generalize to shapes outside of the training distribution by evaluating our model and others on noisy synthetic data composed exclusively of unseen shapes.
-S4.2  Linear Interpolation in Junction Space
-Surprisingly, we note that our network learns a spatially smooth manifold of junctions in its hidden state.
-Figure S4 visualizes some properties of its learned embedding of junction space.
-We generate equally-spaced samples $\boldsymbol{\gamma}_{i}\in\mathbb{R}^{D_{\boldsymbol{\gamma}}}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to 00 and then to a particular $\boldsymbol{\gamma}_{b}$.
-We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center.
-For visual comparison, we show an analogous interpolation from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$ by using a linear geometric interpolation in junction space ${\cal G}$, the expressions for which are in the supplement.
+S4.1  Qualitative Results for Natural Images
 Fig.
-S4: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$.
-Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$.
-The embedding has learned to be smooth and have an intuitive zero.
-S4.3  Evolution of the Outputs
-Figure S5 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement.
-Specifically, we visualize the result of slicing $b_{k}(x;\mathbf{g})$, which we then apply a non-linearity to in order to amplify the less prominent boundaries.
-We see that early iterations are exploratory and unstructured, and that later iterations reach agreement.
-Fig.
-S5: Evolution of boundaries during iterations, in reading order.
-Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
-S4.4  Qualitative Results for Natural Images
-Fig.
-S6: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [19] and Pidinet [23] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [3], and Field of Junctions [25] with patch size $11$.
-In Figures S7 and S6, we show how the model behaves on noiseless natural images that contain texture and recognizable objects.
-In particular, Figure S6 emphasizes how the boundary maps produced by our model qualitatively differ from those of many classical bottom-up edge-detectors and also from those of learned, end-to-end models that have been trained to match human annotations.
-The figure compares our output to that from Canny [3], Field of Junctions [25] with a patch size of $11$, Pidinet [23], and EDTER [19].
-The latter two methods are trained on human annotated data, whereas the former two methods, like our model, are not.
+S3: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [18] and Pidinet [21] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [2], and the field of junctions (FoJ) [23] with patch size $11$.
+In Figures S4 and S3, we show how the model behaves on noiseless natural images that contain texture and recognizable objects.
+In particular, Figure S3 emphasizes how the boundary maps produced by our model qualitatively differ from other methods.
+Here, in addition to showing the results those of classical bottom-up edge-detectors, we include results learned, end-to-end models that have been trained to match human annotations as another point of reference.
+It is important to remember that these methods are trained to identify boundary structures that are defined by semantic variations, whereas our method and other low-level methods divide regions based on variations in color.
+Figure  S3 compares our output to that from Canny [2], the field of junctions (FoJ) [23] with a patch size of $11$, Pidinet [21], and EDTER [18], the latter two being networks trained on human annotated data.
 (Note that inputs for all models besides EDTER [18] were $300\times 400$.
@@
 It does not include mechanisms for grouping based on local texture statistics, nor based on non-local shape and appearance patterns that have semantic meaning to humans.
-Compared to the bottom-up methods of Canny [3] and Field of Junctions [25], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of windowing field $\mathbf{p}[k]$.
-In contrast, the Field of Junctions and Canny both operate at a single pre-determined choice of local size, so they tend to oversegment some places while undersegmenting others.
+Compared to the bottom-up methods of Canny [2] and the field of junctions [23], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of windowing field $\mathbf{p}[k]$.
+In contrast, the the field of junctions and Canny both operate at a single pre-determined choice of local size, so they tend to oversegment some places while undersegmenting others.
 Fig.
-S7: Qualitative behavior of our model on noiseless natural images.
+S4: Qualitative behavior of our model on noiseless natural images.
 From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}[n]$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
@@
 Figure S6 provides additional comparisons for a sample of varying-noise images from the ELD dataset [24].
+We include results from other low level methods and as a point of comparison, the outputs of several methods trained to parse semantic boundaries.
 When detecting boundaries at low signal-to-noise ratios, it is difficult to accurately discern finer structures as the noise level increases.
-Some algorithms, such as Field of Junctions [25], have tunable parameters such as patch-size that provide control over the level of detection.
+Some algorithms, such as the field of junctions (FoJ) [23], have tunable parameters such as patch-size that provide control over the level of detection.
 A small patchsize allows recovering fine structures in lower noise situations, but it causes many false positive boundaries at high noise levels.
@@
 We also note that only the low-level models resolve this level of detail in the first place: The models trained on human annotations—UAED [27], EDTER, HED, Pidinet, and Structured Forests—miss the double contour entirely, estimating instead a single thick curve.
-We emphasize again that a user can adjust the behavior of Canny and Field of Junctions by tuning their local size parameters, either the filter size for Canny or the patchsize for Field of Junctions.
+We emphasize again that a user can adjust the behavior of Canny and the field of junctions by tuning their local size parameters, either the filter size for Canny or the patchsize for the field of junctions.
 Increasing the local size improves their resilience to noise but reduces their spatial precision.
@@
 Fig.
-S8: Visualization of our model’s output for low-light images captured by an iPhone XS.
+S5: Visualization of our model’s output for low-light images captured by an iPhone XS.
 From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
 Fig.
-S9: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise.
-We compare to end-to-end models that are trained to match human annotations (UAED [28], EDTER [19], HED [27], Pidinet [23], and Structured Edges (SE) [7]) in addition to low-level models that are not (Canny [3], gPb [14], and Field of Junctions (FoJ-17) [25]) with patch size 17.
+S6: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise.
+We compare to end-to-end models that are trained to match human annotations (UAED [27], EDTER [18], HED [25], Pidinet [21], and Structured Edges (SE) [6]) in addition to low-level models that are not (Canny [2], gPb [13], and the field of junctions (FoJ-17) [23]) with patch size 17.
 Fig. S7:
@@
 Fig.
-S10: (cont.) Additional qualitative comparisons between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, using crops from the ELD dataset under increasing amounts of photographic noise, including very high levels of noise.
+S7: (cont.) Additional qualitative comparisons between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, using crops from the ELD dataset under increasing amounts of photographic noise, including very high levels of noise.
 S6 Additional Uses of Our Model
@@
 Fig.
-S11: Using our model for depth completion in RGBD images.
+S8: Using our model for depth completion in RGBD images.
 Left: Input RGB channels.